### v2

# FALCON: Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN

Abstract
Recently, large language models (LLMs) have achieved significant progress in automated code generation. Despite their strong instruction-following capabilities, these models frequently struggled to align with user intent in the coding scenario. In particular, they were hampered by datasets that lacked diversity and failed to address specialized tasks or edge cases. Furthermore, challenges in supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) led to failures in generating precise, human-intent-aligned code. To tackle these challenges and improve the code generation performance for automated programming systems, we propose Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN (i.e., FALCON). FALCON is structured into two hierarchical levels, from the global level, long-term memory improves code quality by retaining and applying learned knowledge, while from the local level, short-term memory allows for the incorporation of immediate feedback from compilers and AI systems. Additionally, we introduce meta-reinforcement learning with feedback rewards to solve the global-local bi-level optimization problem and enhance the model’s adaptability across diverse code generation tasks. Extensive experiments are conducted and it is found that our technique achieves state-of-the-art performance, leading other reinforcement learning methods by more than 4.5 percentage points on the MBPP benchmark and 6.1 percentage points on the Humaneval benchmark. The open-sourced code is publicly available at https://anonymous.4open.science/r/FALCON-BFE0/README.md.
1 Introduction
The development of Large Language Models (LLMs) has significantly advanced automated code generation (Zan et al., 2023). Models like CodeLLaMA (Roziere et al., 2023) and DeepSeek-Coder (Guo et al., 2024), tailored for code-centric tasks, have demonstrated outstanding performance across programming challenges. While LLMs excel in instruction-following through tuning (Jiang et al., 2024), they often misalign with user intent, making feedback-based adjustments critical. For example, InstructGPT (Ouyang et al., 2022) leverages reinforcement learning with human feedback (RLHF), and CodeRL (Le et al., 2022) uses compilation feedback to refine model performance. Similarly, CompCoder (Wang et al., 2022) enhances code compilability with compiler feedback, and RLTF (Liu et al., 2023a) offers fine-grained feedback on compiler errors. However, current RL frameworks generate compilation errors and overlook non-differentiable features (e.g. coding style) that affect the performance significantly (Jiang et al., 2024). To address these challenges, we propose a reinforcement learning system combining long-term and short-term memory feedback. From the global level, long-term memory tracks trends over time for higher-quality code retrieval, while from the local level, short-term memory captures recent errors and immediate feedback. The main contributions of this paper are as follows:
  • Short-Term and Long-Term Memory for Reinforcement Learning: We propose a dual-memory approach for reinforcement learning in code generation, where short-term memory enables real-time corrections and long-term memory accumulates knowledge from past runs to improve code quality and reduce repetitive mistakes.
  • Non-Differentiable Code Features into Feedback Loops: Our approach addresses the limitation of current RL frameworks by integrating non-differentiable code features like style, readability, and best practices into the feedback loop, ensuring the generated code is both functionally sound and aligned with real-world programming standards.
  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to efficiently generalize across diverse programming tasks, adapt quickly to new environments, and handle a wide range of coding challenges with fewer training iterations.
2 Related Works
2.1 Pre-trained Models for Code Generation
In recent years, pre-trained language models have made significant progress in the field of code generation. Trained on large-scale code corpora, these models have demonstrated powerful code generation capabilities. For example, CodeBERT Feng et al. (2020), a model based on an encoder-only architecture, has shown impressive results. With the advent of in-context learning methods, decoder-only Transformer models have become the dominant technology for language modeling Vaswani et al. (2017). Several models, such as CodeGPT Lu et al. (2021), CodeGeeX Zheng et al. (2024), and DeepSeek-Coder Guo et al. (2024), use Causal Language Modelling (CLM) pretraining, while others like CodeT5 Wang et al. (2021) and AlphaCode Li et al. (2022) utilize an encoder-decoder architecture. Additionally, models like CodeFusion Singh et al. (2023) leverage diffusion-based techniques for code generation. These pre-trained models exhibit great potential in code generation tasks, achieving notable improvements in accuracy through various architectures and training strategies. However, they still face challenges in ensuring the syntactical and functional correctness of the generated code.
2.2 Reinforcement Learning on code
Reinforcement learning (RL) is a method that learns optimal strategies through reward signals from interacting with the environment Fujimoto et al. (2019). It excels in sequence generation tasks, such as enhancing performance in translation and summarization models by improving BLEU and ROUGE scores Ahn et al. (2019). Unlike traditional natural language processing tasks, code generation is more complex. Beyond syntactical correctness, the generated code must be functionally accurate, meaning it should compile and perform as expected in various scenarios. Passing unit tests is important for verifying correctness, safety, and precision, but not sufficient unless the tests are comprehensive and well-designed. Unit tests must cover diverse use cases and edge cases to ensure the code adheres to standards and meets requirements. For instance, the RL-based code generation fine-tuning framework, CodeRL Le et al. (2022), guides fine-tuning by integrating unit test signals with reinforcement learning. PPOCoder Shojaee et al. (2023) improves on CodeRL by employing Proximal Policy Optimization to refine the approach, while RLTF Liu et al. (2023a) incorporates specific error information from the code and multi-granularity feedback alongside an online framework for model training. StepCoder Dou et al. (2024) enhances the code generation process through compiler feedback and segmental optimization. However, current RL-based methods still face some limitations. Primarily, these methods lack detailed feedback on specific types and distributions of programming errors, which is crucial for identifying patterns, understanding root causes, and creating targeted interventions. This makes addressing recurring issues difficult. Additionally, the corrective mechanisms are not robust enough, often failing to provide timely, specific guidance, hindering effective learning. Additionally, the diversity of input-output examples in existing benchmarks is limited, restricting the model’s ability to adapt and generalize to different or unseen problems Liu et al. (2023a).
3 Problem Setting
We aim to enhance the automated code generation capabilities of Large Language Models (LLMs) by addressing key challenges in accuracy, diversity, error correction, and code quality. Formally, given a high-level specification $D$, the task is to generate a sequence of code tokens $W=\{w_{1},w_{2},\ldots,w_{T}\}$ that maximizes the conditional probability $P(W|D,\theta)$, where $\theta$ represents the model parameters. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\log P(W|D,\theta)\right]$. Challenges. Effective code generation is impeded by several factors. Accuracy requires minimizing syntactical or logical errors in $W$ to ensure correct functionality. Diversity of input-output examples $\mathcal{X}$ is often limited, restricting the model’s ability to generalize across varied programming tasks. Efficient error correction mechanisms are necessary to identify and rectify errors in $W$, ensuring robust performance. Maintaining high code quality, which encompasses adherence to coding standards, style guidelines, and managing code complexity $C(W)$, remains a persistent challenge. Additionally, adaptability refers to the model’s capacity to adapt to new tasks and incorporate feedback for continuous improvement, which is constrained without robust memory mechanisms. FALCON Framework. To address these challenges, we propose the FALCON framework. The framework leverages both a long-term memory buffer $\mathcal{M}_{\text{long}}=\{(D_{i},W_{i},T_{i},F_{i})\}_{i=1}^{N}$ and a short-term memory buffer $\mathcal{M}_{\text{short}}=\{(D_{j},W_{j},T_{j},F_{j})\}_{j=1}^{M}$ to utilize diverse feedback, enabling fine-tuning of $\theta$. Formally, FALCON seeks to optimize $\theta$ by maximizing a composite reward function $R(W,F)=\alpha T(W)+\beta S(W)+\gamma C(W)+\delta E(W)$, where $T(W)$, $S(W)$, $C(W)$, and $E(W)$ represent unit test, code style, complexity, and error feedback, respectively. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\mathbb{E}_{W\sim P(W|D,\theta)}\left[R(W,F)\right]\right]$. Assumptions. The effectiveness of FALCON is based on the following assumptions: 1. Exchangeability of the dataset $\mathcal{D}=\{D_{i}\}_{i=1}^{N}$, implying that the order of tasks does not influence model performance; 2. Independence of feedback signals $F$ given the generated code $W$; 3. Sufficient memory capacity in both $\mathcal{M}_{\text{long}}$ and $\mathcal{M}_{\text{short}}$ to store relevant interactions without significant data loss; and 4. Feedback efficacy, ensuring that integrated feedback mechanisms provide meaningful and actionable information to guide the optimization of $\theta$. Optimization Objective. Given a set of tasks $\mathcal{T}=\{D_{i}\}_{i=1}^{N}$, the goal is to learn optimal parameters $\theta^{*}$ that maximize the expected composite reward across all tasks, formulated as $\theta^{*}=\arg\max_{\theta}\frac{1}{N}\sum_{i=1}^{N}R(W_{i},F_{i})$, where $W_{i}\sim P(W|D_{i},\theta)$ is the generated code for task $D_{i}$, and $F_{i}$ represents the aggregated feedback. The formulation ensures continuous improvement by leveraging both historical and recent feedback to enhance code generation quality.
4 Methodology
In this section, we explore the FALCON framework, which integrates comprehensive unit testing with reinforcement learning, supported by both long-term and short-term memory buffers. During the code generation process, the system stores task descriptions, generated code, and various feedback (e.g., compilation results, code style, and complexity) in the long-term memory buffer. By retrieving this information, the model references high-quality code, avoids past mistakes, and ensures adherence to required standards. After generating the code, a judge model evaluates it and calculates rewards based on the feedback, which are then used to update the model’s parameters through reinforcement learning. All generated code and feedback are stored for future reference and optimization. The combination of long-term and short-term memory feedback in the FALCON framework allows the model to not only learn from a wide range of historical data but also adapt quickly to new tasks based on recent performance. The overall framework is illustrated in Figure 1.
Figure 1: The overview of the FALCON framework. Detailed feedback process (see Appendix Figures 5 and 6 for feedback cases). Reinforcement learning updates using feedback to compute RL loss and optimize the model (see Figure 2 for the detailed meta-reinforcement learning framework).
4.1 Task Definition: Code Generation by LLMs
The code generation task involves the automated creation of computer code $W$ from a high-level specification of the desired behavior $D$. The goal of this process is to enhance the likelihood of generating code that effectively solves a given problem, which is represented as a conditional probability $P(W|D)$. This task can be modeled as a sequence-to-sequence problem, aiming to maximize the conditional probability of generating the correct output given the input and the parameters of the LLM model, mathematically represented as:
$\max_{\theta}P(W|D,\theta)=\max_{\theta}\prod_{t=1}^{T}p(w_{t}|D,\theta,w_{1:t-1})$ (1)
where $\theta$ denotes the parameters of the LLM model. The LLM model is employed to learn the conditional distribution and is trained on a set of input-output pairs. During the training phase, the parameters $\theta$ are optimized to increase the likelihood of generating accurate outputs for a given input.
4.2 MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
We propose the MRLF algorithm 1, which involves random task sampling from the task distribution during training. Previous works have indicated that different data sampling strategies have varying impacts on model information extraction Zhang et al. (2024). Consequently, we implement both long and short memory sequences. The long memory strategy stores the solutions generated for each problem and the compiler feedback results, whereas the short memory sequence selects the latest samples and unit test feedback from each current iteration. To address repetitive runtime and compilation errors, the long memory strategy categorizes and stores various errors by their types and records the corresponding error lines, enabling fine-grained reward allocation.
Algorithm 1 MRLF Algorithm
1:Task distribution $\mathcal{T}$
2:Updated model parameters $\theta$
3:Initialize $LMB$, $SMB$, $\theta$
4:Populate $LMB$, $SMB$ via few-shot demonstrations
5:repeat
6:     Sample batch $\{T_{i}\}$ from $\mathcal{T}$
7:     for each task $T_{i}$ do
8:         Code Generation: Generate code $\hat{W}$ and test results; record in $SMB$
9:         Adaptation: Set $\theta_{i}=\theta$; sample mini-batch from $LMB$, $SMB$
[…]
11:         Update $\theta_{i}$: $\theta_{i}\leftarrow\theta_{i}-\alpha\nabla_{\theta_{i}}L_{\text{inner}}(\theta_{i})$
12:         Evaluation: Assess $\theta_{i}$ on $T_{i}$; document outcomes
13:     end for
14:     Meta-Update: $L_{\text{meta}}=\sum_{i=1}^{N}L_{\text{inner}}(\theta_{i})$; refine $\theta$: $\theta\leftarrow\theta-\beta\nabla_{\theta}L_{\text{meta}}$
15:until Convergence
16:return $\theta$
We initialize two memory buffers (LMB, SMB) and model parameters $\theta$, then randomly sample tasks from the task distribution, executing steps for each one. First, the corresponding code is generated and tested, with results stored in SMB. During adaptation, the algorithm extracts experimental data from LMB and SMB to aid rapid adaptation to the current task. The inner loss function $L_{\text{inner}}$, incorporating factors like accuracy, complexity, style, and negative examples, is optimized using policy gradient. In evaluation, the optimized parameters are tested, and results are stored for future use. Finally, the losses from all tasks are aggregated to compute the meta-loss $L_{\text{meta}}$, used to update global parameters $\theta$. Algorithm 1 summarizes the framework. For inner loop optimization, our method explores the target space by combining unit test feedback, code complexity, and style norms, using the generated code $\hat{w}$ to construct the reinforcement learning loss function as shown below:
$L_{r1}=-\sum_{t=S_{\text{fine}}}^{E_{\text{fine}}}R_{\text{fine}}(\hat{w}_{t})\log p(\hat{w}_{t}|D,\theta,\hat{w}_{1:t-1})$ (2)
where $R_{\text{fine}}(\ast)$ represents the reward coefficient, and $S_{\text{fine}}$ and $E_{\text{fine}}$ denote the start and end positions of the code snippet, respectively. These values are determined based on different types of feedback. To stabilize the training process, we adopt the supervised learning loss $L_{sl}$ by minimizing the cross-entropy loss as shown below:
$L_{sl}=-\log P(w\mid D,\theta)=-\sum_{t=1}^{T}\log p(w_{t}\mid D,\theta,w_{1:t-1})$ (3)
Figure 2: The framework integrates meta-reinforcement learning with both long-term and short-term memory feedback. Long-term memory retrieves historical tasks, while short-term memory provides AI and compiler feedback to refine code generation.
As depicted in Figure 2, we employ a meta-reinforcement learning framework to optimize code generation by integrating both long- and short-term memories for enhanced adaptability. From a global perspective, Long-Term Memory $D_{\text{long}}$ stores historical tasks, generated codes, and feedback to provide valuable context. From a local perspective, Short-Term Memory $D_{\text{short}}$ focuses on recent feedback to enable real-time adjustments. This approach leverages the MAML framework (Finn et al., 2017) for efficient task adaptation with minimal updates.
Short-Term Memory Adaptation. Short-Term Memory is utilized to adapt the model locally by adjusting its parameters based on recent feedback. For each task $\mathcal{T}_{i}$, the inner loop optimization updates the parameters:
$\theta_{i}^{\prime}=\theta_{i}-\alpha\nabla_{\theta_{i}}\mathcal{L}_{\mathcal{T}_{i}}(\theta_{i})$ (4)
where $\alpha$ is the learning rate and $\mathcal{L}_{\mathcal{T}_{i}}$ is the task-specific loss function.
Global Optimization. The outer loop performs global optimization of the meta-learning parameters $\theta_{\text{meta}}$ by aggregating feedback across multiple tasks:
$\theta_{\text{meta}}=\theta_{\text{meta}}-\beta\nabla_{\theta_{\text{meta}}}\sum_{i}\mathcal{L}_{\mathcal{T}_{i}}(\theta_{i}^{\prime})$ (5)
where $\beta$ is the meta-learning rate. This ensures better generalization across tasks.
Final Optimization. The overall framework combines short- and long-term memory feedbacks with meta-reinforcement learning to achieve coordinated optimization for both global generalization and local task adaptation:
$\theta_{\text{final}}=\text{Optimize}\left(\theta_{\text{meta}},\theta,\{\theta_{i}^{\prime}\}\right)$ (6)
4.3 Long-Term Memory Feedback
Retrieving information from long-term memory helps improve code quality. We use the FAISS framework (Douze et al., 2024) to retrieve relevant historical code, feedback, and evaluation scores. Task descriptions and feedback are transformed into embedding vectors and then indexed. During code generation, a query vector from the current task retrieves the top-k most similar historical data to guide the process and avoid past errors. The prompt template is provided in the appendix. Consider a set of historical data $\mathcal{D}={(t_{i},f_{i},e_{i})}_{i=1}^{n}$, where $t_{i}$ represents the task description, $f_{i}$ is the corresponding feedback, and $e_{i}$ is the evaluation score. We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\bm{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS. During the code generation phase, the current task description $\bm{t}_{\text{current}}$ and feedback $\bm{f}$ are transformed into a query vector $\bm{q}=\phi(\bm{t}_{\text{current}})$. We compute the similarity between the query vector $\bm{q}$ and the historical vectors $\bm{v}_{i}$ using cosine similarity $\text{cos}(\bm{q},\bm{v}_{i})$, and retrieve the top-$k$ most similar historical tasks. The retrieval process can be represented as:
$\{(t_{i_{1}},f_{i_{1}},e_{i_{1}}),\dots,(t_{i_{k}},f_{i_{k}},e_{i_{k}})\}=\text{Top-}k\left(\bm{v}_{i}\mid i=1,2,\dots,n\right)$ (7)
By referencing these most relevant historical tasks and feedbacks, the system can guide the current code generation process with past mistakes avoided and ultimate code quality improved.
4.4 Short-Term Memory Feedback
During the reinforcement learning phase, we utilize the generated code $\hat{w}$ to construct the reinforcement learning loss function as follows:
$L_{r1}=-\sum_{t=S_{\text{fine}}}^{E_{\text{fine}}}R_{\text{fine}}(\hat{w}_{t})\log p(\hat{w}_{t}|D,\theta,\hat{w}_{1:t-1})$ (8)
where $R_{\text{fine}}(\ast)$ represents the reward coefficient, and $S_{\text{fine}}$ and $E_{\text{fine}}$ denote the start and end positions of the code snippet, respectively. These values are determined based on different types of feedback. Compiler Feedback. For compiler feedback, we adopt the same settings as CodeRL:
$R_{\text{coarse}}(\hat{W})=\left\{\begin{array}[]{ll}\phantom{-}1.0,&\text{if }FB(\hat{W})\text{ is pass}\\ -0.3,&\text{if }FB(\hat{W})\text{ is failure}\\ -0.6,&\text{if }FB(\hat{W})\text{ is runtime error}\\ -1.0,&\text{if }FB(\hat{W})\text{ is syntax error}\end{array}\right.$ (9)
$S_{\text{coarse}}=0,\quad E_{\text{coarse}}=T$
where $R_{\text{coarse}}$ is based on compiler feedback with the start and end positions set to 0 and $T$.
Adaptive Feedback. To enhance the model’s efficiency in handling various programming tasks, we devise a mechanism that dynamically adjusts rewards based on the proportion of passes to failures in unit tests. This strategy encourages the model not only to pass unit tests but also to learn from failures, thereby improving its problem-solving capabilities. The reward is calculated as:
$R_{\text{error}}(\hat{W})=-0.3+1.3\times\frac{N_{\text{pass}}}{N_{\text{pass}}+N_{\text{fail}}}$ (10)
Coding Style Feedback. To further enhance the quality of the generated code, we employ AI Feedback to optimize coding style. An evaluation model scores the generated code based on adherence to the expected coding style standards. The scoring system ranges from -1 to 2, and these evaluation scores are directly used as reward signals in the reinforcement learning process to guide the model toward producing higher-quality code. The coding style assessment template is provided in Table 8.
Complexity Feedback. Just like with coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores. The complexity assessment template is provided in Table 9.
Long-term Error Type Feedback. Introducing a reward mechanism that combines short-term memory recall error rate, current test error, and long-term memory recall of past task performance enables the model to dynamically adjust its rewards based on past error patterns, adapt to various error types and feedback, and ultimately enhance its generalization ability:
$R_{\text{negative}}=-\sum_{\text{error}}N_{\text{error}}\times P_{\text{error}}$ (11)
where $N_{\text{error}}$ represents the short-term memory recall, the frequency of each error type in the generated code by the model. This is the immediate feedback for the current task. $P_{\text{error}}$ represents the long-term memory recall, the proportion of each error type. This is the model’s performance statistics on long-term tasks, providing a cumulative history of various error types. By correct rewarding, the model can reduce the occurrence of these errors and thereby enhance the accuracy and quality of the generated code.
5 Experiment
5.1 Quantitative Evaluation on APPS
To ensure a fair comparison, we use the CodeT5 770M model as our baseline. Our benchmarks include the latest advancements that integrate reinforcement learning (RL) with large language models (LLMs), particularly CodeRL, PPOCoder, and RLTF. For evaluation, we apply the same benchmarks and settings used in these previous works. As shown in Table 1 for the experimental results, our FALCON approach delivers additional performance improvements and surpasses other RL-based methods, indicating that RL with appropriate feedback can effectively improve the model output space and thereby enhance the quality of code generation. In particular, our method achieves the highest pass@1 rates of 8.60%, 2.56%, and 1.25% in the Introductory, Interview, and Competition categories, respectively.
Table 1: Quantitative evaluation on the APPS benchmark. “Intro”: introductory, “Inter”: interview, “Comp”: competition-level tasks. To ensure a fair comparison, we apply these RL-based methods, including PPOCoder, CodeRL, and RLTF, using the same base model, CodeT5, as a backbone. We also compare with models that have a larger number of parameters.
Method Size pass@1 pass@5
Intro Inter Comp All Intro Inter Comp All
Codex 12B 4.14 0.14 0.02 0.92 9.65 0.51 0.09 2.25
GPT-Neo 2.7B 3.90 0.57 0 1.12 5.50 0.80 0 1.58
CodeT5 base 770M 3.85 0.58 0.02 1.05 8.52 1.53 0.25 2.82
PPOCoder 770M 4.06 0.79 0.15 1.32 9.97 2.06 0.70 3.37
CodeRL 770M 7.08 1.86 0.75 2.69 16.37 4.95 2.84 6.81
RLTF 770M 8.40 2.28 1.10 3.27 18.60 5.57 3.70 7.87
Ours 770M 8.60 2.56 1.24 3.50 19.75 5.85 3.57 8.17
5.2 Quantitative Evaluation on HumanEval and MBPP
To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model, trained with our method on our custom dataset, using the well-established MBPP and HumanEval benchmarks. We also compare these results against other reinforcement learning methods, such as PPOCoder and RLTF. The experimental results are illustrated in Table 2.
Table 2: The results of pass@1 on the MBPP and HumanEval benchmarks.
Model Humaneval MBPP
DeepSeek-Coder-Instruct 73.8 74.9
PPOCoder 76.8 76.2
RLTF 76.8 75.9
Ours 82.9 80.7
Compared to other reinforcement learning methods, our method consistently achieves the best performance on both the HumanEval and MBPP benchmarks. The significant advantage of our method can be attributed to its diversified feedback mechanism. Unlike other methods that may focus on a single metric, our method continuously optimizes the model’s generation capability through multi-dimensional feedback. This approach demonstrates a strong ability to enhance the generation of correct code and proves particularly effective in complex tasks.
5.3 Quantitative Evaluation on CODAL-Bench
In addition to evaluating the functional correctness of the code, we adopt CODAL-Bench, a rigorous and comprehensive benchmark for LLM consistency in coding preferences to validate the effectiveness of short-term memory feedback. DeepSeek-Coder-Instruct-6.7B model is used and the results are illustrated in Figure 3. It is found that there is a noticeable improvement in various coding preferences, particularly in Code Complexity and Coding Style after implementing the FALCON framework. This observation is attributed to the inclusion of feedback on these aspects in short-term memory. However, the improvement in Instruction Following is not as significant.
Figure 3: Quantitative evaluation on CODAL-Bench
5.4 Quantitative Evaluation on SciCode
To validate the general-purpose task capabilities of our framework, we also select the SciCode benchmark, which covers challenging research-level coding problems across natural sciences, including mathematics, physics, chemistry, and biology. SciCode decomposes the main problems into several subproblems, making it a particularly rigorous benchmark of a model’s coding capabilities. Even the most advanced models nowadays, such as Claude 3.5-Sonnet and ChatGPT-4.0, can only solve 1.5% and 4.6% of the main problems, respectively. Although Deepseek-Coder-6.7B-instruct initially demonstrates a low task pass rate on this benchmark, we observe improvements on the subproblems after applying our framework due to the utilization of long-term memory mechanisms.
Table 3: The results of pass@1 on SciCode benchmarks with and without FALCON.
Model Size Method Subproblem Main Problem
CodeLlama 70B - 10.4 0
CodeLlama 7B w/o 0.4 0
CodeLlama 7B w 3.5 0
DeepSeek-Coder 6.7B w/o 5.2 0
DeepSeek-Coder 6.7B w 8.3 0
5.5 Quantitative Evaluation on Agentbench
To further evaluate the performance of our framework, we conduct a systematic assessment on AgentBench Liu et al. (2023b), focusing specifically on long-term memory capabilities. Since our primary focus is on code generation tasks, we select three environments within AgentBench: Operating System (OS), Database (DB), and Knowledge Graph (KG). In this evaluation, we compare proprietary models (such as GPT-4 and GPT-3.5) with open-source models (such as Codellama and Qwen 2.5) Liang et al. (2024). The results reveal that the models optimized through our framework exhibit significant improvements, particularly in the OS environment with an increase of 4.8 percentage points. The experimental results are illustrated in Table 4.
Table 4: Test set results of AGENTBENCH.
Model Size VER OS DB KG
GPT-3.5-turbo - 0613 31.6 15.7 25.9
GPT-4 - 0613 42.4 32.0 58.8
Tinyllama 1.1B - 2.8 0 0
Codellama 7B instruct 9.7 2.7 0
Qwen 7B chat 12.5 13.0 7.0
Agentlm 7B chat 14.6 33 9.0
Deepseek-Coder 6.7B instruct 17.4 23.3 6.8
DeepSeek-Coder (FALCON) 6.7B instruct 22.2 26.7 9.0
5.6 Ablation Studies
The Influence of Models. To validate the scalability and robustness of our framework, we conduct experiments with the larger model, DeepSeek-Coder-Instruct-6.7B, to further evaluate its performance. Notably, the improvements in introductory-level tasks are significant, which can be attributed to the use of long-term memory that enhances the quality of generated data and further unlocks the model’s potential. The results are illustrated in Table 5.
Table 5: Different large language models as the backbone
Model Size Method Intro Inter Comp All
CodeT5 770M w/o 3.85 0.58 0.02 1.12
CodeT5 770M w 8.60 2.56 1.25 3.50
DeepSeek-Coder 6.7B w/o 16.70 7.20 2.30 8.12
DeepSeek-Coder 6.7B w 22.40 8.52 3.70 10.33
The Influence of Different Feedbacks on Coding Preferences. As shown in Figure 4, ablation experiments are also conducted to validate the effectiveness of the feedback that we introduce for coding preferences. It is found that incorporating targeted feedback enhances the model’s performance concerning its respective coding preferences. Notably, the optimization aiming at increasing code complexity achieves the best results. Although there are some improvements in coding style and instruction following, it is worth noting that the enhancement in code instruction following is not significant. We leave it as a topic for future investigation.
Figure 4: Effect of different feedbacks on coding preferences
The Influence of Memory. To validate the impact of long-term and short-term memories on code generation capabilities, we conduct ablation experiments using CodeT5 as the base model and test it on the APPS dataset. As shown in Table 6, the experimental results indicate that both long- and short-term memory feedback enhances the model’s code generation performance effectively, while the short-term memory feedback demonstrates a more significant improvement. This improvement can be attributed to the effective reward design which plays a positive role in fine-tuning the model.
Table 6: Effect of long and short memories on different performance metrics
Long Memory Short Memory Intro Inter Comp All
- - 3.85 0.58 0.02 1.12
- 4.14 0.74 0.02 1.28
- 7.20 1.86 0.70 2.70
8.60 2.56 1.25 3.50
6 Conclusions and Future Work
In this work, we propose FALCON, a novel RL-based feedback He et al. (2024) framework that enhances automated code generation by integrating long-term and short-term memory feedbacks within a meta-reinforcement learning strategy to better capture human intent like RL based preference learning. Long-term memory retains past interactions to improve code quality and reduce repetitive mistakes, while short-term memory enables immediate adjustments based on recent feedback from compilers and AI systems. This dual-memory approach addresses limitations in existing models that struggle to align code generation with user intent, especially in specialized tasks or edge cases. By incorporating non-differentiable code features like style and complexity into the feedback loop, FALCON ensures that the generated code is not only functionally correct but also adheres to real-world programming standards. Extensive evaluations on benchmarks including APPS, HumanEval, and CODAL-Bench demonstrate that FALCON outperforms existing RL-based methods, achieving higher functional correctness and better coding style adherence. In future work, we aim to expand FALCON’s capabilities by incorporating a broader diversity of programming languages and tackling more complex code generation challenges.
Appendix A Example
We provided an example of code generation with long-term memory. When generating code without long-term memory, it often results in repetitive ValueError issues. By incorporating long-term memory to retrieve the most relevant code blocks and embedding them as context during generation, the quality of the generated code can be significantly improved. We also provided an instance of using long-term memory in SCICode. In direct generation, there were certain logical issues within the code, and the coding style lumped all formulas together. However, after employing long-term memory retrieval for assistance, the code was segmented appropriately and the logic was correctly implemented.
Figure 5: An example of code generation with long-term memory incorporated.
Figure 6: An example of code generation with long-term memory incorporated.
Appendix B Feedback Error Correlation Analysis
We performed an analysis of different feedback types and their associated error categories. We conducted experiments using CodeT5 as the base model on APPS with individual feedback and collected the occurrence frequency of various sub-errors. The results are shown in the Figure 7. The experimental results indicate that compiler feedback significantly reduces Syntax Errors and Index Errors. However, it also slightly increases the occurrence of Value Errors. This can be attributed to the corrective nature of compiler feedback on errors. Other feedback types, such as Coding Style, Instruction Following, and Code Complexity Feedback, can reduce Syntax Errors compared to having no feedback. However, their reduction is not as significant as that achieved by compiler feedback. Instruction Following Feedback specifically shows some improvement in reducing Value Errors, indicating an enhancement in instruction adherence.
Figure 7: Frequency of sub-errors generated by different feedback types.
Appendix C Prompts
We have compiled relevant prompt templates for code generation based on long-term memory retrieval and AI feedback.
Table 7: Code generation template
Instruction:
Please write a Python function based on the task description, referencing the historical context for inspiration. Ensure that the generated code follows the provided requirements and avoids the listed errors.
Task: [instruction]
Context of relevant code:
- Historical Task: [Brief description of a similar task]
- Code: [Code snippet]
- Style Score: [Style score]
- Efficiency Score: [Efficiency score]
- Additional Feedback: [Additional comments or issues]
Requirements:
1. Ensure the generated code adheres to best practices for Python, including proper naming conventions, consistent formatting, and coding standards.
2. Optimize the code for performance, avoiding unnecessary recursion or nested loops.
3. Use built-in or efficient library functions whenever applicable to improve both readability and performance.
Avoid the following errors:
- [Historical error] – Avoid structural or logical issues found in previous code snippets.
Output: Ensure your response is in the format of ”’python”’.
Table 8: Coding style assessment template.
Coding Style Assessment
Evaluate the coding style of provided code segments. Assess how well the code adheres to the best practices of the language, focusing on readability, maintainability, and efficiency in line with the language’s idiomatic style.
Reward Scale: Rate outputs on a scale of -1 to 2
-1. Poor Adherence: The code significantly deviates from standard practices, showing poor readability, maintainability, and efficiency.
0. Basic Adherence: The code makes some effort to follow language conventions but lacks consistency in readability, maintainability, or efficiency.
1. Good Adherence: The code generally follows standards, demonstrating adequate readability, maintainability, and efficiency, though with room for improvement.
2. Excellent Adherence: The code exemplifies best practices, with high readability, maintainability, and efficiency, fully adhering to idiomatic conventions.
Table 9: Complexity assessment template.
Complexity Assessment
Evaluate the solutions and code provided by the assistant based on their complexity. Assess how well the code manages complexity while achieving the desired outcomes.
Reward Scale: Rate outputs on a scale of -1 to 2
-1. Overly Complex: The code is unnecessarily complicated, with a high level of complexity that makes it hard to understand or maintain.
0. Acceptable Complexity: The code has a reasonable level of complexity, but there may be opportunities to simplify further.
1. Moderately Simple: The code is simple and well-organized, with minimal complexity and clear logic.
2. Optimal Simplicity: The code exemplifies the best practices in minimizing complexity while ensuring functionality.
Table 10: Instruction following assessment template.
Instruction Following Assessment
Evaluate the assistant’s fidelity to provided instructions. Assess how accurately the assistant’s responses align with user directives, noting any deviations and their justification.
Evaluation Criteria
Precision in Following Instructions: Does the assistant adhere to the specifics of the provided instructions?
Justification for Deviations: If deviations occur, are they justified by critical necessity or explicit user request?
Alignment with User Directives: How well do the assistant’s responses match the user’s specified needs and expectations?
Necessity of Deviations: Are any deviations from instructions made only in situations deemed absolutely necessary or upon direct user request?
Reward Scale: Rate outputs on a scale of -1 to 2
-1. Non-Compliant: The assistant frequently deviates from instructions without necessity or user consent.
0. Acceptable: The assistant shows some adherence to instructions but deviates without strong justification.
1. Compliant with Justified Deviations: The assistant generally follows instructions, with deviations occurring but justified by necessity or user request.
2. Fully Compliant: The assistant follows instructions closely, with minimal deviations, all of which are well justified.
Appendix D Error Category
Due to the differences in languages accepted by Compiler Feedback during unit tests for various language tasks, we have standardized the definition of sub-errors in Compiler Feedback. The table 11 12 13 below outlines our specifications for Python, C, and Java.
Table 11: Common Python errors with categories.
Sub-error Description Category
Syntax Error Code contains syntax errors that cause the compilation to fail. Syntax Error
Indentation Error Wrong indentation format. Syntax Error
Index Error Index operation is out of bounds. Runtime Error
Type Error An operation or function was applied to an object of an inappropriate type. Runtime Error
Value Error An operation or function received an argument with the correct type but with an inappropriate value. Runtime Error
EOF Error The input() function encountered an end-of-file condition (EOF) without reading any data. Runtime Error
Timeout Error Code execution time exceeds time limit. Runtime Error
Name Error A local or global name is not defined. Runtime Error
Key Error A mapping (dictionary) key is not found in the set of existing keys. Runtime Error
Import Error The imported package is not found. Runtime Error
ZeroDivision Error The second argument of a division or modulo operation is zero. Runtime Error
Recursion Error Code execution recursive operation exceeds the maximum limit. Runtime Error
Table 12: Common C Language errors with categories.
Sub-error Description Category
Segmentation Fault Accessing memory that the program doesn’t have permission to access. Runtime Error
Null Pointer Dereference Attempting to dereference a pointer that is NULL. Runtime Error
Buffer Overflow Writing data outside the allocated buffer memory. Runtime Error
Memory Leak Dynamically allocated memory not being freed. Runtime Error
Syntax Error A syntax mistake in the code, such as a missing semicolon. Syntax Error
Type Mismatch Assigning a value of one type to a variable of another type. Syntax Error
Uninitialized Variable Using a variable before it has been initialized. Runtime Error
Undefined Behavior Code that can exhibit unpredictable behavior depending on compiler or runtime environment. Runtime Error
Division by Zero Attempting to divide a number by zero. Runtime Error
Infinite Loop A loop that never terminates due to incorrect logic. Runtime Error
Table 13: Common Java Language errors with categories.
Sub-error Description Category
NullPointerException Attempting to access an object with a null reference. Runtime Error
ArrayIndexOutOfBounds- Exception Accessing an array index that is out of bounds. Runtime Error
ClassCastException Casting an object to a subclass it is not an instance of. Runtime Error
NumberFormatException Attempting to convert a string to a number, but the string doesn’t have the appropriate format. Runtime Error
StackOverflowError Recursive method calls that exceed the stack size. Runtime Error
Syntax Error Any mistake in the code structure such as missing braces or semicolons. Syntax Error
ClassNotFoundException The Java class is not found at runtime. Runtime Error
IllegalArgumentException A method has been passed an illegal or inappropriate argument. Runtime Error
ArithmeticException Division by zero or other illegal arithmetic operations. Runtime Error
UnsupportedOperation- Exception When a requested operation is not supported. Runtime Error
Appendix E Supplementary Experiments
E.1 Quantitative Evaluation on APPS
We utilized a larger model, DeepSeek-Coder-Instruct-6.7B, as the backbone to compare popular and widely used advanced reinforcement learning methods, further validating the effectiveness and necessity of our approach. The specific results indicate that our method achieved scores of 44.7, 14.4, and 3.2 in the categories of "Introductory" "Interview" and "Competition" respectively. This demonstrates our method’s exceptional capability in exploring the output space of larger models, highlighting its versatility.
Table 14: Results of pass@1 on APPS.
Models Size Introductory Interview Competition Overall
Base Models
CodeLlama 13B 14.0 5.0 1.0 6.0
CodeLlama-Python 13B 25.3 8.2 2.4 10.5
DeepSeek-Coder-Base 6.7B 11.4 8.5 2.8 7.9
Supervised Fine-tuned Models
StarCoder 15.6B 6.2 2.9 0.3 3.1
CodeLlama-Instruct 13B 30.1 11.8 3.2 13.7
WizardCoder-Python-V1.0 13B 35.2 10.9 4.1 14.4
DeepSeek-Coder-Instruct 6.7B 40.2 13.7 2.7 16.8
Reinforcement Learning-based Models (DeepSeek-Coder-Instruct-6.7B as the backbone)
PPOCoder 6.7B 41.2 14.0 2.2 17.1
CodeRL 6.7B 40.8 13.7 2.7 16.9
RLTF 6.7B 42.4 13.3 3.1 17.1
FALCON(Ours) 6.7B 44.7 14.4 3.2 18.2
E.2 Quantitative Evaluation on LeetCode
To further validate our approach in solving real-world programming problems, we utilized the LeetCode Contest benchmark. This benchmark comprises 180 of the latest LeetCode contest problems collected from July 2023 to January 2024. Each problem is accompanied by 100 test cases, and the data format is consistent with human-eval. The results indicate that our approach can enhance the code generation capabilities of models in solving real-world programming problems. Compared to other RL-based methods, our approach achieved greater improvements and can rival models with significantly larger parameters.
Table 15: Results of pass@1 on LeetCode.
Models Size Easy(45) Medium(91) Hard(44) Overall(180)
GPT-3.5-Turbo - 46.7 15.4 15.9 23.3
WizardCoder-Python-V1.0 13B 17.8 1.1 0 5
CodeLlama-Instruct 34B 24.4 4.4 4.5 9.4
DeepSeek-Coder-Instruct 6.7B 44.4 12.1 9.1 19.4
Reinforcement Learning-based Models
PPOCoder 6.7B 48.9 13.1 9.1 21.1
CodeRL 6.7B 55.6 14.2 9.1 23.3
RLTF 6.7B 57.8 13.1 11.3 23.9
FALCON(Ours) 6.7B 60.0 15.4 11.3 25.6

### v3

@@
 Extensive experiments are conducted and it is found that our technique achieves state-of-the-art performance, leading other reinforcement learning methods by more than 4.5 percentage points on the MBPP benchmark and 6.1 percentage points on the Humaneval benchmark.
-The open-sourced code is publicly available at https://anonymous.4open.science/r/FALCON-BFE0/README.md.
-1 Introduction
-The development of Large Language Models (LLMs) has significantly advanced automated code generation (Zan et al., 2023).
-Models like CodeLLaMA (Roziere et al., 2023) and DeepSeek-Coder (Guo et al., 2024), tailored for code-centric tasks, have demonstrated outstanding performance across programming challenges.
-While LLMs excel in instruction-following through tuning (Jiang et al., 2024), they often misalign with user intent, making feedback-based adjustments critical.
-For example, InstructGPT (Ouyang et al., 2022) leverages reinforcement learning with human feedback (RLHF), and CodeRL (Le et al., 2022) uses compilation feedback to refine model performance.
-Similarly, CompCoder (Wang et al., 2022) enhances code compilability with compiler feedback, and RLTF (Liu et al., 2023a) offers fine-grained feedback on compiler errors.
+The open-sourced code is publicly available at https://anonymous.4open.science/r/FALCON-3B64/README.md.
+Index Terms:
+Code generation, Reinforcement Learning, Diverse Feedback.
+I Introduction
+The development of Large Language Models (LLMs) has significantly advanced automated code generation [1].
+Models like CodeLLaMA [2] and DeepSeek-Coder [3], tailored for code-centric tasks, have demonstrated outstanding performance across programming challenges.
+While LLMs excel in instruction-following through tuning [4], they often misalign with user intent, making feedback-based adjustments critical.
+For example, InstructGPT [5] leverages reinforcement learning with human feedback (RLHF), and CodeRL [6] uses compilation feedback to refine model performance.
+Similarly, CompCoder [7] enhances code compilability with compiler feedback, and RLTF [8] offers fine-grained feedback on compiler errors.
 However, current RL frameworks generate compilation errors and overlook non-differentiable features (e.g.
-coding style) that affect the performance significantly (Jiang et al., 2024).
+coding style) that affect the performance significantly [4].
 To address these challenges, we propose a reinforcement learning system combining long-term and short-term memory feedback.
@@
   • Non-Differentiable Code Features into Feedback Loops: Our approach addresses the limitation of current RL frameworks by integrating non-differentiable code features like style, readability, and best practices into the feedback loop, ensuring the generated code is both functionally sound and aligned with real-world programming standards.
-  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to efficiently generalize across diverse programming tasks, adapt quickly to new environments, and handle a wide range of coding challenges with fewer training iterations.
-2 Related Works
-2.1 Pre-trained Models for Code Generation
+  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to generalize across diverse programming tasks efficiently, adapt to new environments quickly, and handle a wide range of coding challenges with fewer training iterations.
+II Related Works
+II-A Pre-trained Models for Code Generation
 In recent years, pre-trained language models have made significant progress in the field of code generation.
 Trained on large-scale code corpora, these models have demonstrated powerful code generation capabilities.
-For example, CodeBERT Feng et al.
-(2020), a model based on an encoder-only architecture, has shown impressive results.
-With the advent of in-context learning methods, decoder-only Transformer models have become the dominant technology for language modeling Vaswani et al.
-(2017).
-Several models, such as CodeGPT Lu et al.
-(2021), CodeGeeX Zheng et al.
-(2024), and DeepSeek-Coder Guo et al.
-(2024), use Causal Language Modelling (CLM) pretraining, while others like CodeT5 Wang et al.
-(2021) and AlphaCode Li et al.
-(2022) utilize an encoder-decoder architecture.
-Additionally, models like CodeFusion Singh et al.
-(2023) leverage diffusion-based techniques for code generation.
+For example, CodeBERT [9], a model based on an encoder-only architecture, has shown impressive results.
+With the advent of in-context learning methods, decoder-only Transformer models have become the dominant technology for language modeling [10].
+Several models, such as CodeGPT [11], CodeGeeX [12], and DeepSeek-Coder [3], use Causal Language Modelling (CLM) pretraining, while others like CodeT5 [13] and AlphaCode [14] utilize an encoder-decoder architecture.
+Additionally, models like CodeFusion [15] leverage diffusion-based techniques for code generation.
 These pre-trained models exhibit great potential in code generation tasks, achieving notable improvements in accuracy through various architectures and training strategies.
 However, they still face challenges in ensuring the syntactical and functional correctness of the generated code.
-2.2 Reinforcement Learning on code
-Reinforcement learning (RL) is a method that learns optimal strategies through reward signals from interacting with the environment Fujimoto et al.
-(2019).
-It excels in sequence generation tasks, such as enhancing performance in translation and summarization models by improving BLEU and ROUGE scores Ahn et al.
-(2019).
+II-B Reinforcement Learning on code
+Reinforcement learning (RL) is a method that learns optimal strategies through reward signals from interacting with the environment [16].
+It excels in sequence generation tasks, such as enhancing performance in translation and summarization models by improving BLEU and ROUGE scores [17].
 Unlike traditional natural language processing tasks, code generation is more complex.
@@
 Unit tests must cover diverse use cases and edge cases to ensure the code adheres to standards and meets requirements.
-For instance, the RL-based code generation fine-tuning framework, CodeRL Le et al.
-(2022), guides fine-tuning by integrating unit test signals with reinforcement learning.
-PPOCoder Shojaee et al.
-(2023) improves on CodeRL by employing Proximal Policy Optimization to refine the approach, while RLTF Liu et al.
-(2023a) incorporates specific error information from the code and multi-granularity feedback alongside an online framework for model training.
-StepCoder Dou et al.
-(2024) enhances the code generation process through compiler feedback and segmental optimization.
+For instance, the RL-based code generation fine-tuning framework, CodeRL [6], guides fine-tuning by integrating unit test signals with reinforcement learning.
+PPOCoder [18] improves on CodeRL by employing Proximal Policy Optimization to refine the approach, while RLTF [8] incorporates specific error information from the code and multi-granularity feedback alongside an online framework for model training.
+StepCoder [19] enhances the code generation process through compiler feedback and segmental optimization.
 However, current RL-based methods still face some limitations.
@@
 Additionally, the corrective mechanisms are not robust enough, often failing to provide timely, specific guidance, hindering effective learning.
-Additionally, the diversity of input-output examples in existing benchmarks is limited, restricting the model’s ability to adapt and generalize to different or unseen problems Liu et al.
-(2023a).
-3 Problem Setting
+Additionally, the diversity of input-output examples in existing benchmarks is limited, restricting the model’s ability to adapt and generalize to different or unseen problems [8].
+III Problem Setting
 We aim to enhance the automated code generation capabilities of Large Language Models (LLMs) by addressing key challenges in accuracy, diversity, error correction, and code quality.
@@
 The formulation ensures continuous improvement by leveraging both historical and recent feedback to enhance code generation quality.
-4 Methodology
+IV Methodology
 In this section, we explore the FALCON framework, which integrates comprehensive unit testing with reinforcement learning, supported by both long-term and short-term memory buffers.
@@
 Reinforcement learning updates using feedback to compute RL loss and optimize the model (see Figure 2 for the detailed meta-reinforcement learning framework).
-4.1 Task Definition: Code Generation by LLMs
+IV-A Task Definition: Code Generation by LLMs
 The code generation task involves the automated creation of computer code $W$ from a high-level specification of the desired behavior $D$.
@@
 During the training phase, the parameters $\theta$ are optimized to increase the likelihood of generating accurate outputs for a given input.
-4.2 MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
+IV-B MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
 We propose the MRLF algorithm 1, which involves random task sampling from the task distribution during training.
-Previous works have indicated that different data sampling strategies have varying impacts on model information extraction Zhang et al.
-(2024).
+Previous works have indicated that different data sampling strategies have varying impacts on model information extraction [20].
 Consequently, we implement both long and short memory sequences.
@@
 Algorithm 1 MRLF Algorithm
-1:Task distribution $\mathcal{T}$
-2:Updated model parameters $\theta$
-3:Initialize $LMB$, $SMB$, $\theta$
-4:Populate $LMB$, $SMB$ via few-shot demonstrations
-5:repeat
-6:     Sample batch $\{T_{i}\}$ from $\mathcal{T}$
-7:     for each task $T_{i}$ do
-8:         Code Generation: Generate code $\hat{W}$ and test results; record in $SMB$
-9:         Adaptation: Set $\theta_{i}=\theta$; sample mini-batch from $LMB$, $SMB$
+Task distribution $\mathcal{T}$
+Updated model parameters $\theta$
+Initialize $LMB$, $SMB$, $\theta$
+Populate $LMB$, $SMB$ via few-shot demonstrations
+repeat
+     Sample batch $\{T_{i}\}$ from $\mathcal{T}$
+     for each task $T_{i}$ do
+         Code Generation: Generate code $\hat{W}$ and test results; record in $SMB$
+         Adaptation: Set $\theta_{i}=\theta$; sample mini-batch from $LMB$, $SMB$
 […]
-11:         Update $\theta_{i}$: $\theta_{i}\leftarrow\theta_{i}-\alpha\nabla_{\theta_{i}}L_{\text{inner}}(\theta_{i})$
-12:         Evaluation: Assess $\theta_{i}$ on $T_{i}$; document outcomes
-13:     end for
-14:     Meta-Update: $L_{\text{meta}}=\sum_{i=1}^{N}L_{\text{inner}}(\theta_{i})$; refine $\theta$: $\theta\leftarrow\theta-\beta\nabla_{\theta}L_{\text{meta}}$
-15:until Convergence
-16:return $\theta$
+         Update $\theta_{i}$: $\theta_{i}\leftarrow\theta_{i}-\alpha\nabla_{\theta_{i}}L_{\text{inner}}(\theta_{i})$
+         Evaluation: Assess $\theta_{i}$ on $T_{i}$; document outcomes
+     end for
+     Meta-Update: $L_{\text{meta}}=\sum_{i=1}^{N}L_{\text{inner}}(\theta_{i})$; refine $\theta$: $\theta\leftarrow\theta-\beta\nabla_{\theta}L_{\text{meta}}$
+until Convergence
+return $\theta$
 We initialize two memory buffers (LMB, SMB) and model parameters $\theta$, then randomly sample tasks from the task distribution, executing steps for each one.
@@
 For inner loop optimization, our method explores the target space by combining unit test feedback, code complexity, and style norms, using the generated code $\hat{w}$ to construct the reinforcement learning loss function as shown below:
-$L_{r1}=-\sum_{t=S_{\text{fine}}}^{E_{\text{fine}}}R_{\text{fine}}(\hat{w}_{t})\log p(\hat{w}_{t}|D,\theta,\hat{w}_{1:t-1})$ (2)
+$L_{rl}=-\sum_{t=S_{\text{fine}}}^{E_{\text{fine}}}R_{\text{fine}}(\hat{w}_{t})\log p(\hat{w}_{t}|D,\theta,\hat{w}_{1:t-1})$ (2)
 where $R_{\text{fine}}(\ast)$ represents the reward coefficient, and $S_{\text{fine}}$ and $E_{\text{fine}}$ denote the start and end positions of the code snippet, respectively.
@@
 $L_{sl}=-\log P(w\mid D,\theta)=-\sum_{t=1}^{T}\log p(w_{t}\mid D,\theta,w_{1:t-1})$ (3)
-Figure 2: The framework integrates meta-reinforcement learning with both long-term and short-term memory feedback.
-Long-term memory retrieves historical tasks, while short-term memory provides AI and compiler feedback to refine code generation.
 As depicted in Figure 2, we employ a meta-reinforcement learning framework to optimize code generation by integrating both long- and short-term memories for enhanced adaptability.
@@
 From a local perspective, Short-Term Memory $D_{\text{short}}$ focuses on recent feedback to enable real-time adjustments.
-This approach leverages the MAML framework (Finn et al., 2017) for efficient task adaptation with minimal updates.
+This approach leverages the MAML framework [21] for efficient task adaptation with minimal updates.
+Figure 2: The framework integrates meta-reinforcement learning with both long-term and short-term memory feedback.
+Long-term memory retrieves historical tasks, while short-term memory provides AI and compiler feedback to refine code generation.
 Short-Term Memory Adaptation.
@@
 $\theta_{\text{final}}=\text{Optimize}\left(\theta_{\text{meta}},\theta,\{\theta_{i}^{\prime}\}\right)$ (6)
-4.3 Long-Term Memory Feedback
-Retrieving information from long-term memory helps improve code quality.
-We use the FAISS framework (Douze et al., 2024) to retrieve relevant historical code, feedback, and evaluation scores.
+IV-C Long-Term Memory Feedback
+Given that retrieving information from long-term memory helps improve code quality, we use the FAISS framework [22] to retrieve relevant historical code, feedback, and evaluation scores.
 Task descriptions and feedback are transformed into embedding vectors and then indexed.
@@
 Consider a set of historical data $\mathcal{D}={(t_{i},f_{i},e_{i})}_{i=1}^{n}$, where $t_{i}$ represents the task description, $f_{i}$ is the corresponding feedback, and $e_{i}$ is the evaluation score.
-We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\bm{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS.
-During the code generation phase, the current task description $\bm{t}_{\text{current}}$ and feedback $\bm{f}$ are transformed into a query vector $\bm{q}=\phi(\bm{t}_{\text{current}})$.
-We compute the similarity between the query vector $\bm{q}$ and the historical vectors $\bm{v}_{i}$ using cosine similarity $\text{cos}(\bm{q},\bm{v}_{i})$, and retrieve the top-$k$ most similar historical tasks.
+We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\boldsymbol{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS.
+During the code generation phase, the current task description $\boldsymbol{t}_{\text{current}}$ and feedback $\boldsymbol{f}$ are transformed into a query vector $\boldsymbol{q}=\phi(\boldsymbol{t}_{\text{current}})$.
+We compute the similarity between the query vector $\boldsymbol{q}$ and the historical vectors $\boldsymbol{v}_{i}$ using cosine similarity $\text{cos}(\boldsymbol{q},\boldsymbol{v}_{i})$, and retrieve the top-$k$ most similar historical tasks.
 The retrieval process can be represented as:
-$\{(t_{i_{1}},f_{i_{1}},e_{i_{1}}),\dots,(t_{i_{k}},f_{i_{k}},e_{i_{k}})\}=\text{Top-}k\left(\bm{v}_{i}\mid i=1,2,\dots,n\right)$ (7)
+$\{(t_{i_{1}},f_{i_{1}},e_{i_{1}}),\dots,(t_{i_{k}},f_{i_{k}},e_{i_{k}})\}=\text{Top-}k\left(\boldsymbol{v}_{i}\mid i=1,2,\dots,n\right)$ (7)
 By referencing these most relevant historical tasks and feedbacks, the system can guide the current code generation process with past mistakes avoided and ultimate code quality improved.
-4.4 Short-Term Memory Feedback
+IV-D Short-Term Memory Feedback
 During the reinforcement learning phase, we utilize the generated code $\hat{w}$ to construct the reinforcement learning loss function as follows:
@@
 These values are determined based on different types of feedback.
-Compiler Feedback.
-For compiler feedback, we adopt the same settings as CodeRL:
+Compiler Feedback. For compiler feedback, we adopt the same settings as CodeRL:
 $R_{\text{coarse}}(\hat{W})=\left\{\begin{array}[]{ll}\phantom{-}1.0,&\text{if }FB(\hat{W})\text{ is pass}\\ -0.3,&\text{if }FB(\hat{W})\text{ is failure}\\ -0.6,&\text{if }FB(\hat{W})\text{ is runtime error}\\ -1.0,&\text{if }FB(\hat{W})\text{ is syntax error}\end{array}\right.$ (9)
-$S_{\text{coarse}}=0,\quad E_{\text{coarse}}=T$
-where $R_{\text{coarse}}$ is based on compiler feedback with the start and end positions set to 0 and $T$.
+where $R_{\text{coarse}}$ is based on compiler feedback with the start position $S_{\text{coarse}}=0$ and end position $E_{\text{coarse}}=T$.
 Adaptive Feedback.
@@
 The scoring system ranges from -1 to 2, and these evaluation scores are directly used as reward signals in the reinforcement learning process to guide the model toward producing higher-quality code.
-The coding style assessment template is provided in Table 8.
-Complexity Feedback. Just like with coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores. The complexity assessment template is provided in Table 9.
+The coding style assessment template is provided in Table X.
+Complexity Feedback. Similar to the coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores. The complexity assessment template is provided in Table XI.
 Long-term Error Type Feedback.
-Introducing a reward mechanism that combines short-term memory recall error rate, current test error, and long-term memory recall of past task performance enables the model to dynamically adjust its rewards based on past error patterns, adapt to various error types and feedback, and ultimately enhance its generalization ability:
+Introducing a reward mechanism that combines short-term memory recall error rate, current test error, and long-term memory feedback of past task performance enables the model to dynamically adjust its rewards based on past error patterns, adapt to various error types and feedback, and ultimately enhance its generalization ability:
 $R_{\text{negative}}=-\sum_{\text{error}}N_{\text{error}}\times P_{\text{error}}$ (11)
-where $N_{\text{error}}$ represents the short-term memory recall, the frequency of each error type in the generated code by the model.
-This is the immediate feedback for the current task.
-$P_{\text{error}}$ represents the long-term memory recall, the proportion of each error type.
+where $N_{\text{error}}$ represents the frequency of each error type in the generated code, which is the short-term memory feedback and the immediate feedback for the current task.
+$P_{\text{error}}$ represents the proportion of each error type, which is the long-term memory feedback.
 This is the model’s performance statistics on long-term tasks, providing a cumulative history of various error types.
 By correct rewarding, the model can reduce the occurrence of these errors and thereby enhance the accuracy and quality of the generated code.
-5 Experiment
-5.1 Quantitative Evaluation on APPS
-To ensure a fair comparison, we use the CodeT5 770M model as our baseline.
-Our benchmarks include the latest advancements that integrate reinforcement learning (RL) with large language models (LLMs), particularly CodeRL, PPOCoder, and RLTF.
-For evaluation, we apply the same benchmarks and settings used in these previous works.
-As shown in Table 1 for the experimental results, our FALCON approach delivers additional performance improvements and surpasses other RL-based methods, indicating that RL with appropriate feedback can effectively improve the model output space and thereby enhance the quality of code generation.
-In particular, our method achieves the highest pass@1 rates of 8.60%, 2.56%, and 1.25% in the Introductory, Interview, and Competition categories, respectively.
-Table 1: Quantitative evaluation on the APPS benchmark.
+TABLE I: Quantitative evaluation on the APPS benchmark.
 “Intro”: introductory, “Inter”: interview, “Comp”: competition-level tasks.
@@
 Ours 770M 8.60 2.56 1.24 3.50 19.75 5.85 3.57 8.17
-5.2 Quantitative Evaluation on HumanEval and MBPP
-To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model, trained with our method on our custom dataset, using the well-established MBPP and HumanEval benchmarks.
+V Experiment
+V-A Quantitative Evaluation on APPS
+To ensure a fair comparison, we use the CodeT5 770M model as our baseline.
+Our benchmarks include the latest advancements that integrate reinforcement learning (RL) with large language models (LLMs), particularly CodeRL, PPOCoder, and RLTF.
+For evaluation, we apply the same benchmarks and settings used in these previous works.
+As shown in Table I for the experimental results, our FALCON approach delivers additional performance improvements and surpasses other RL-based methods, indicating that RL with appropriate feedback can effectively improve the model output space and thereby enhance the quality of code generation.
+In particular, our method achieves the highest pass@1 rates of 8.60%, 2.56%, and 1.25% in the Introductory, Interview, and Competition categories, respectively.
+V-B Quantitative Evaluation on HumanEval and MBPP
+To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model trained with our method on APPS, using the well-established MBPP and HumanEval benchmarks.
 We also compare these results against other reinforcement learning methods, such as PPOCoder and RLTF.
-The experimental results are illustrated in Table 2.
-Table 2: The results of pass@1 on the MBPP and HumanEval benchmarks.
+The experimental results are illustrated in Table II.
+TABLE II: The results of pass@1 on the MBPP and HumanEval benchmarks.
 Model Humaneval MBPP
@@
 This approach demonstrates a strong ability to enhance the generation of correct code and proves particularly effective in complex tasks.
-5.3 Quantitative Evaluation on CODAL-Bench
-In addition to evaluating the functional correctness of the code, we adopt CODAL-Bench, a rigorous and comprehensive benchmark for LLM consistency in coding preferences to validate the effectiveness of short-term memory feedback.
-DeepSeek-Coder-Instruct-6.7B model is used and the results are illustrated in Figure 3.
-It is found that there is a noticeable improvement in various coding preferences, particularly in Code Complexity and Coding Style after implementing the FALCON framework.
-This observation is attributed to the inclusion of feedback on these aspects in short-term memory.
-However, the improvement in Instruction Following is not as significant.
-Figure 3: Quantitative evaluation on CODAL-Bench
-5.4 Quantitative Evaluation on SciCode
-To validate the general-purpose task capabilities of our framework, we also select the SciCode benchmark, which covers challenging research-level coding problems across natural sciences, including mathematics, physics, chemistry, and biology.
-SciCode decomposes the main problems into several subproblems, making it a particularly rigorous benchmark of a model’s coding capabilities.
-Even the most advanced models nowadays, such as Claude 3.5-Sonnet and ChatGPT-4.0, can only solve 1.5% and 4.6% of the main problems, respectively.
-Although Deepseek-Coder-6.7B-instruct initially demonstrates a low task pass rate on this benchmark, we observe improvements on the subproblems after applying our framework due to the utilization of long-term memory mechanisms.
-Table 3: The results of pass@1 on SciCode benchmarks with and without FALCON.
-Model Size Method Subproblem Main Problem
-CodeLlama 70B - 10.4 0
-CodeLlama 7B w/o 0.4 0
-CodeLlama 7B w 3.5 0
-DeepSeek-Coder 6.7B w/o 5.2 0
-DeepSeek-Coder 6.7B w 8.3 0
-5.5 Quantitative Evaluation on Agentbench
-To further evaluate the performance of our framework, we conduct a systematic assessment on AgentBench Liu et al.
-(2023b), focusing specifically on long-term memory capabilities.
-Since our primary focus is on code generation tasks, we select three environments within AgentBench: Operating System (OS), Database (DB), and Knowledge Graph (KG).
-In this evaluation, we compare proprietary models (such as GPT-4 and GPT-3.5) with open-source models (such as Codellama and Qwen 2.5) Liang et al.
-(2024).
-The results reveal that the models optimized through our framework exhibit significant improvements, particularly in the OS environment with an increase of 4.8 percentage points.
-The experimental results are illustrated in Table 4.
-Table 4: Test set results of AGENTBENCH.
+TABLE III: Test set results of AGENTBENCH.
 Model Size VER OS DB KG
@@
 DeepSeek-Coder (FALCON) 6.7B instruct 22.2 26.7 9.0
-5.6 Ablation Studies
+V-C Quantitative Evaluation on CODAL-Bench
+In addition to evaluating the functional correctness of the code, we adopt CODAL-Bench, a rigorous and comprehensive benchmark for LLM consistency in coding preferences to validate the effectiveness of short-term memory feedback.
+DeepSeek-Coder-Instruct-6.7B model is used and the results are illustrated in Figure 3.
+It is found that there is a noticeable improvement in various coding preferences, particularly in Code Complexity and Coding Style after implementing the FALCON framework.
+This observation is attributed to the inclusion of feedback on these aspects in short-term memory.
+However, the improvement in Instruction Following is not as significant.
+Figure 3: Quantitative evaluation on CODAL-Bench
+V-D Quantitative Evaluation on SciCode
+To validate the general-purpose task capabilities of our framework, we also select the SciCode benchmark, which covers challenging research-level coding problems across natural sciences, including mathematics, physics, chemistry, and biology.
+SciCode decomposes the main problems into several subproblems, making it a particularly rigorous benchmark of a model’s coding capabilities.
+Even the most advanced models nowadays, such as Claude 3.5-Sonnet and ChatGPT-4.0, can only solve 1.5% and 4.6% of the main problems, respectively.
+Although Deepseek-Coder-6.7B-instruct initially demonstrates a low task pass rate on this benchmark, we observe improvements on the subproblems after applying our framework due to the utilization of long-term memory mechanisms.
+TABLE IV: The results of pass@1 on SciCode benchmarks with and without FALCON.
+Model Size Method Subproblem Main Problem
+CodeLlama 70B - 10.4 0
+CodeLlama 7B w/o 0.4 0
+CodeLlama 7B w 3.5 0
+DeepSeek-Coder 6.7B w/o 5.2 0
+DeepSeek-Coder 6.7B w 8.3 0
+V-E Quantitative Evaluation on Agentbench
+To further evaluate the performance of our framework, we conduct a systematic assessment on AgentBench [23], focusing specifically on long-term memory capabilities.
+Since our primary focus is on code generation tasks, we select three environments within AgentBench: Operating System (OS), Database (DB), and Knowledge Graph (KG).
+In this evaluation, we compare proprietary models (such as GPT-4 and GPT-3.5) with open-source models (such as Codellama and Qwen 2.5) [24].
+The results reveal that the models optimized through our framework exhibit significant improvements, particularly in the OS environment with an increase of 4.8 percentage points.
+The experimental results are illustrated in Table III.
+V-F Ablation Studies
 The Influence of Models.
-To validate the scalability and robustness of our framework, we conduct experiments with the larger model, DeepSeek-Coder-Instruct-6.7B, to further evaluate its performance.
+To validate the scalability and robustness of our framework, we conduct experiments with DeepSeek-Coder-Instruct-6.7B.
 Notably, the improvements in introductory-level tasks are significant, which can be attributed to the use of long-term memory that enhances the quality of generated data and further unlocks the model’s potential.
-The results are illustrated in Table 5.
-Table 5: Different large language models as the backbone
+The results are illustrated in Table V.
+TABLE V: Different large language models as the backbone
 Model Size Method Intro Inter Comp All
@@
 To validate the impact of long-term and short-term memories on code generation capabilities, we conduct ablation experiments using CodeT5 as the base model and test it on the APPS dataset.
-As shown in Table 6, the experimental results indicate that both long- and short-term memory feedback enhances the model’s code generation performance effectively, while the short-term memory feedback demonstrates a more significant improvement.
-This improvement can be attributed to the effective reward design which plays a positive role in fine-tuning the model.
-Table 6: Effect of long and short memories on different performance metrics
+As shown in Table VI, the experimental results indicate that both long- and short-term memories enhance the model’s code generation performance effectively, while the short-term memory feedback demonstrates a more significant improvement attributed to the reward design which plays an important role in fine-tuning the model.
+TABLE VI: Effect of long and short memories on different performance metrics
 Long Memory Short Memory Intro Inter Comp All
@@
 8.60 2.56 1.25 3.50
-6 Conclusions and Future Work
-In this work, we propose FALCON, a novel RL-based feedback He et al.
-(2024) framework that enhances automated code generation by integrating long-term and short-term memory feedbacks within a meta-reinforcement learning strategy to better capture human intent like RL based preference learning.
+VI Conclusions and Future Work
+In this work, we propose FALCON, a novel RL-based feedback framework that enhances automatic code generation by integrating long-term and short-term memory feedback within a meta-reinforcement learning strategy to better capture human intent like RL based preference learning.
 Long-term memory retains past interactions to improve code quality and reduce repetitive mistakes, while short-term memory enables immediate adjustments based on recent feedback from compilers and AI systems.
@@
 We have compiled relevant prompt templates for code generation based on long-term memory retrieval and AI feedback.
-Table 7: Code generation template
+TABLE VII: Code generation template
 Instruction:
@@
 Output: Ensure your response is in the format of ”’python”’.
-Table 8: Coding style assessment template.
-Coding Style Assessment
-Evaluate the coding style of provided code segments.
-Assess how well the code adheres to the best practices of the language, focusing on readability, maintainability, and efficiency in line with the language’s idiomatic style.
-Reward Scale: Rate outputs on a scale of -1 to 2
--1. Poor Adherence: The code significantly deviates from standard practices, showing poor readability, maintainability, and efficiency.
-0. Basic Adherence: The code makes some effort to follow language conventions but lacks consistency in readability, maintainability, or efficiency.
-1. Good Adherence: The code generally follows standards, demonstrating adequate readability, maintainability, and efficiency, though with room for improvement.
-2. Excellent Adherence: The code exemplifies best practices, with high readability, maintainability, and efficiency, fully adhering to idiomatic conventions.
-Table 9: Complexity assessment template.
-Complexity Assessment
-Evaluate the solutions and code provided by the assistant based on their complexity. Assess how well the code manages complexity while achieving the desired outcomes.
-Reward Scale: Rate outputs on a scale of -1 to 2
--1. Overly Complex: The code is unnecessarily complicated, with a high level of complexity that makes it hard to understand or maintain.
-0. Acceptable Complexity: The code has a reasonable level of complexity, but there may be opportunities to simplify further.
-1. Moderately Simple: The code is simple and well-organized, with minimal complexity and clear logic.
-2. Optimal Simplicity: The code exemplifies the best practices in minimizing complexity while ensuring functionality.
-Table 10: Instruction following assessment template.
-Instruction Following Assessment
-Evaluate the assistant’s fidelity to provided instructions. Assess how accurately the assistant’s responses align with user directives, noting any deviations and their justification.
-Evaluation Criteria
-Precision in Following Instructions: Does the assistant adhere to the specifics of the provided instructions?
-Justification for Deviations: If deviations occur, are they justified by critical necessity or explicit user request?
-Alignment with User Directives: How well do the assistant’s responses match the user’s specified needs and expectations?
-Necessity of Deviations: Are any deviations from instructions made only in situations deemed absolutely necessary or upon direct user request?
-Reward Scale: Rate outputs on a scale of -1 to 2
--1. Non-Compliant: The assistant frequently deviates from instructions without necessity or user consent.
-0. Acceptable: The assistant shows some adherence to instructions but deviates without strong justification.
-1. Compliant with Justified Deviations: The assistant generally follows instructions, with deviations occurring but justified by necessity or user request.
-2. Fully Compliant: The assistant follows instructions closely, with minimal deviations, all of which are well justified.
-Appendix D Error Category
-Due to the differences in languages accepted by Compiler Feedback during unit tests for various language tasks, we have standardized the definition of sub-errors in Compiler Feedback.
-The table 11 12 13 below outlines our specifications for Python, C, and Java.
-Table 11: Common Python errors with categories.
-Sub-error Description Category
-Syntax Error Code contains syntax errors that cause the compilation to fail. Syntax Error
-Indentation Error Wrong indentation format. Syntax Error
-Index Error Index operation is out of bounds. Runtime Error
-Type Error An operation or function was applied to an object of an inappropriate type. Runtime Error
-Value Error An operation or function received an argument with the correct type but with an inappropriate value. Runtime Error
-EOF Error The input() function encountered an end-of-file condition (EOF) without reading any data. Runtime Error
-Timeout Error Code execution time exceeds time limit. Runtime Error
-Name Error A local or global name is not defined. Runtime Error
-Key Error A mapping (dictionary) key is not found in the set of existing keys. Runtime Error
-Import Error The imported package is not found. Runtime Error
-ZeroDivision Error The second argument of a division or modulo operation is zero. Runtime Error
-Recursion Error Code execution recursive operation exceeds the maximum limit. Runtime Error
-Table 12: Common C Language errors with categories.
-Sub-error Description Category
-Segmentation Fault Accessing memory that the program doesn’t have permission to access. Runtime Error
-Null Pointer Dereference Attempting to dereference a pointer that is NULL. Runtime Error
-Buffer Overflow Writing data outside the allocated buffer memory. Runtime Error
-Memory Leak Dynamically allocated memory not being freed. Runtime Error
-Syntax Error A syntax mistake in the code, such as a missing semicolon. Syntax Error
-Type Mismatch Assigning a value of one type to a variable of another type. Syntax Error
-Uninitialized Variable Using a variable before it has been initialized. Runtime Error
-Undefined Behavior Code that can exhibit unpredictable behavior depending on compiler or runtime environment. Runtime Error
-Division by Zero Attempting to divide a number by zero. Runtime Error
-Infinite Loop A loop that never terminates due to incorrect logic. Runtime Error
-Table 13: Common Java Language errors with categories.
-Sub-error Description Category
-NullPointerException Attempting to access an object with a null reference. Runtime Error
-ArrayIndexOutOfBounds- Exception Accessing an array index that is out of bounds. Runtime Error
-ClassCastException Casting an object to a subclass it is not an instance of. Runtime Error
-NumberFormatException Attempting to convert a string to a number, but the string doesn’t have the appropriate format. Runtime Error
-StackOverflowError Recursive method calls that exceed the stack size. Runtime Error
-Syntax Error Any mistake in the code structure such as missing braces or semicolons. Syntax Error
-ClassNotFoundException The Java class is not found at runtime. Runtime Error
-IllegalArgumentException A method has been passed an illegal or inappropriate argument. Runtime Error
-ArithmeticException Division by zero or other illegal arithmetic operations. Runtime Error
-UnsupportedOperation- Exception When a requested operation is not supported. Runtime Error
-Appendix E Supplementary Experiments
-E.1 Quantitative Evaluation on APPS
-We utilized a larger model, DeepSeek-Coder-Instruct-6.7B, as the backbone to compare popular and widely used advanced reinforcement learning methods, further validating the effectiveness and necessity of our approach.
-The specific results indicate that our method achieved scores of 44.7, 14.4, and 3.2 in the categories of "Introductory" "Interview" and "Competition" respectively.
-This demonstrates our method’s exceptional capability in exploring the output space of larger models, highlighting its versatility.
-Table 14: Results of pass@1 on APPS.
+TABLE VIII: Results of pass@1 on APPS.
 Models Size Introductory Interview Competition Overall
@@
 FALCON(Ours) 6.7B 44.7 14.4 3.2 18.2
-E.2 Quantitative Evaluation on LeetCode
-To further validate our approach in solving real-world programming problems, we utilized the LeetCode Contest benchmark.
-This benchmark comprises 180 of the latest LeetCode contest problems collected from July 2023 to January 2024.
-Each problem is accompanied by 100 test cases, and the data format is consistent with human-eval.
-The results indicate that our approach can enhance the code generation capabilities of models in solving real-world programming problems.
-Compared to other RL-based methods, our approach achieved greater improvements and can rival models with significantly larger parameters.
-Table 15: Results of pass@1 on LeetCode.
+TABLE IX: Results of pass@1 on LeetCode.
 Models Size Easy(45) Medium(91) Hard(44) Overall(180)
@@
 FALCON(Ours) 6.7B 60.0 15.4 11.3 25.6
+TABLE X: Coding style assessment template.
+Coding Style Assessment
+Evaluate the coding style of provided code segments.
+Assess how well the code adheres to the best practices of the language, focusing on readability, maintainability, and efficiency in line with the language’s idiomatic style.
+Reward Scale: Rate outputs on a scale of -1 to 2
+-1. Poor Adherence: The code significantly deviates from standard practices, showing poor readability, maintainability, and efficiency.
+0. Basic Adherence: The code makes some effort to follow language conventions but lacks consistency in readability, maintainability, or efficiency.
+1. Good Adherence: The code generally follows standards, demonstrating adequate readability, maintainability, and efficiency, though with room for improvement.
+2. Excellent Adherence: The code exemplifies best practices, with high readability, maintainability, and efficiency, fully adhering to idiomatic conventions.
+TABLE XI: Complexity assessment template.
+Complexity Assessment
+Evaluate the solutions and code provided by the assistant based on their complexity. Assess how well the code manages complexity while achieving the desired outcomes.
+Reward Scale: Rate outputs on a scale of -1 to 2
+-1. Overly Complex: The code is unnecessarily complicated, with a high level of complexity that makes it hard to understand or maintain.
+0. Acceptable Complexity: The code has a reasonable level of complexity, but there may be opportunities to simplify further.
+1. Moderately Simple: The code is simple and well-organized, with minimal complexity and clear logic.
+2. Optimal Simplicity: The code exemplifies the best practices in minimizing complexity while ensuring functionality.
+TABLE XII: Instruction following assessment template.
+Instruction Following Assessment
+Evaluate the assistant’s fidelity to provided instructions. Assess how accurately the assistant’s responses align with user directives, noting any deviations and their justification.
+Evaluation Criteria
+Precision in Following Instructions: Does the assistant adhere to the specifics of the provided instructions?
+Justification for Deviations: If deviations occur, are they justified by critical necessity or explicit user request?
+Alignment with User Directives: How well do the assistant’s responses match the user’s specified needs and expectations?
+Necessity of Deviations: Are any deviations from instructions made only in situations deemed absolutely necessary or upon direct user request?
+Reward Scale: Rate outputs on a scale of -1 to 2
+-1. Non-Compliant: The assistant frequently deviates from instructions without necessity or user consent.
+0. Acceptable: The assistant shows some adherence to instructions but deviates without strong justification.
+1. Compliant with Justified Deviations: The assistant generally follows instructions, with deviations occurring but justified by necessity or user request.
+2. Fully Compliant: The assistant follows instructions closely, with minimal deviations, all of which are well justified.
+Appendix D Error Category
+Due to the differences in languages accepted by Compiler Feedback during unit tests for various language tasks, we have standardized the definition of sub-errors in Compiler Feedback.
+The table XIII XIV XV below outlines our specifications for Python, C, and Java.
+TABLE XIII: Common Python errors with categories.
+Sub-error Description Category
+Syntax Error Code contains syntax errors that cause the compilation to fail. Syntax Error
+Indentation Error Wrong indentation format. Syntax Error
+Index Error Index operation is out of bounds. Runtime Error
+Type Error An operation or function was applied to an object of an inappropriate type. Runtime Error
+Value Error An operation or function received an argument with the correct type but with an inappropriate value. Runtime Error
+EOF Error The input() function encountered an end-of-file condition (EOF) without reading any data. Runtime Error
+Timeout Error Code execution time exceeds time limit. Runtime Error
+Name Error A local or global name is not defined. Runtime Error
+Key Error A mapping (dictionary) key is not found in the set of existing keys. Runtime Error
+Import Error The imported package is not found. Runtime Error
+ZeroDivision Error The second argument of a division or modulo operation is zero. Runtime Error
+Recursion Error Code execution recursive operation exceeds the maximum limit. Runtime Error
+TABLE XIV: Common C Language errors with categories.
+Sub-error Description Category
+Segmentation Fault Accessing memory that the program doesn’t have permission to access. Runtime Error
+Null Pointer Dereference Attempting to dereference a pointer that is NULL. Runtime Error
+Buffer Overflow Writing data outside the allocated buffer memory. Runtime Error
+Memory Leak Dynamically allocated memory not being freed. Runtime Error
+Syntax Error A syntax mistake in the code, such as a missing semicolon. Syntax Error
+Type Mismatch Assigning a value of one type to a variable of another type. Syntax Error
+Uninitialized Variable Using a variable before it has been initialized. Runtime Error
+Undefined Behavior Code that can exhibit unpredictable behavior depending on compiler or runtime environment. Runtime Error
+Division by Zero Attempting to divide a number by zero. Runtime Error
+Infinite Loop A loop that never terminates due to incorrect logic. Runtime Error
+TABLE XV: Common Java Language errors with categories.
+Sub-error Description Category
+NullPointerException Attempting to access an object with a null reference. Runtime Error
+ArrayIndexOutOfBounds- Exception Accessing an array index that is out of bounds. Runtime Error
+ClassCastException Casting an object to a subclass it is not an instance of. Runtime Error
+NumberFormatException Attempting to convert a string to a number, but the string doesn’t have the appropriate format. Runtime Error
+StackOverflowError Recursive method calls that exceed the stack size. Runtime Error
+Syntax Error Any mistake in the code structure such as missing braces or semicolons. Syntax Error
+ClassNotFoundException The Java class is not found at runtime. Runtime Error
+IllegalArgumentException A method has been passed an illegal or inappropriate argument. Runtime Error
+ArithmeticException Division by zero or other illegal arithmetic operations. Runtime Error
+UnsupportedOperation- Exception When a requested operation is not supported. Runtime Error
+Appendix E Supplementary Experiments
+E-A Quantitative Evaluation on APPS
+We utilized a larger model, DeepSeek-Coder-Instruct-6.7B, as the backbone to compare popular and widely used advanced reinforcement learning methods, further validating the effectiveness and necessity of our approach.
+The specific results indicate that our method achieved scores of 44.7, 14.4, and 3.2 in the categories of "Introductory" "Interview" and "Competition" respectively.
+This demonstrates our method’s exceptional capability in exploring the output space of larger models, highlighting its versatility.
+E-B Quantitative Evaluation on LeetCode
+To further validate our approach in solving real-world programming problems, we utilized the LeetCode Contest benchmark.
+This benchmark comprises 180 of the latest LeetCode contest problems collected from July 2023 to January 2024.
+Each problem is accompanied by 100 test cases, and the data format is consistent with human-eval.
+The results indicate that our approach can enhance the code generation capabilities of models in solving real-world programming problems.
+Compared to other RL-based methods, our approach achieved greater improvements and can rival models with significantly larger parameters.

### v5

@@
 To tackle these challenges and improve the code generation performance for automated programming systems, we propose Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN (i.e., FALCON).
-FALCON is structured into two hierarchical levels, from the global level, long-term memory improves code quality by retaining and applying learned knowledge, while from the local level, short-term memory allows for the incorporation of immediate feedback from compilers and AI systems.
-Additionally, we introduce meta-reinforcement learning with feedback rewards to solve the global-local bi-level optimization problem and enhance the model’s adaptability across diverse code generation tasks.
-Extensive experiments are conducted and it is found that our technique achieves state-of-the-art performance, leading other reinforcement learning methods by more than 4.5 percentage points on the MBPP benchmark and 6.1 percentage points on the Humaneval benchmark.
-The open-sourced code is publicly available at https://anonymous.4open.science/r/FALCON-3B64/README.md.
+FALCON leverages long-term memory to retain and apply learned knowledge, short-term memory to incorporate immediate feedback, and meta-reinforcement learning with feedback rewards to address global-local bi-level optimization and enhance adaptability across diverse code generation tasks.Extensive experiments show that FALCON achieves state-of-the-art performance, outperforming other reinforcement learning methods by over 4.5% on MBPP and 6.1% on Humaneval, with the code publicly available.https://anonymous.4open.science/r/FALCON-3B64/README.md.
 Index Terms:
 Code generation, Reinforcement Learning, Diverse Feedback.
+Figure 1: The overview of the FALCON framework.
+Detailed feedback process.
+Reinforcement learning updates using feedback to compute RL loss and optimize the model (see Figure 2 for the detailed meta-reinforcement learning framework).
 I Introduction
@@
 From the global level, long-term memory tracks trends over time for higher-quality code retrieval, while from the local level, short-term memory captures recent errors and immediate feedback.
-The main contributions of this paper are as follows:
+In summary, the main contributions of this paper are as follows:
   • Short-Term and Long-Term Memory for Reinforcement Learning: We propose a dual-memory approach for reinforcement learning in code generation, where short-term memory enables real-time corrections and long-term memory accumulates knowledge from past runs to improve code quality and reduce repetitive mistakes.
   • Non-Differentiable Code Features into Feedback Loops: Our approach addresses the limitation of current RL frameworks by integrating non-differentiable code features like style, readability, and best practices into the feedback loop, ensuring the generated code is both functionally sound and aligned with real-world programming standards.
-  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to generalize across diverse programming tasks efficiently, adapt to new environments quickly, and handle a wide range of coding challenges with fewer training iterations.
+  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to efficiently generalize across diverse programming tasks, adapt quickly to new environments, and handle a wide range of coding challenges with fewer training iterations.
 II Related Works
@@
 With the advent of in-context learning methods, decoder-only Transformer models have become the dominant technology for language modeling [10].
-Several models, such as CodeGPT [11], CodeGeeX [12], and DeepSeek-Coder [3], use Causal Language Modelling (CLM) pretraining, while others like CodeT5 [13] and AlphaCode [14] utilize an encoder-decoder architecture.
+Several models, such as CodeGeeX [11] and DeepSeek-Coder [3], use Causal Language Modelling (CLM) pretraining, while others like CodeT5 [12] utilize an encoder-decoder architecture.
 Additionally, models like CodeFusion [13] leverage diffusion-based techniques for code generation.
@@
 II-B Reinforcement Learning on code
-Reinforcement learning (RL) is a method that learns optimal strategies through reward signals from interacting with the environment [16].
-It excels in sequence generation tasks, such as enhancing performance in translation and summarization models by improving BLEU and ROUGE scores [17].
-Unlike traditional natural language processing tasks, code generation is more complex.
-Beyond syntactical correctness, the generated code must be functionally accurate, meaning it should compile and perform as expected in various scenarios.
-Passing unit tests is important for verifying correctness, safety, and precision, but not sufficient unless the tests are comprehensive and well-designed.
-Unit tests must cover diverse use cases and edge cases to ensure the code adheres to standards and meets requirements.
-For instance, the RL-based code generation fine-tuning framework, CodeRL [6], guides fine-tuning by integrating unit test signals with reinforcement learning.
-PPOCoder [18] improves on CodeRL by employing Proximal Policy Optimization to refine the approach, while RLTF [8] incorporates specific error information from the code and multi-granularity feedback alongside an online framework for model training.
-StepCoder [19] enhances the code generation process through compiler feedback and segmental optimization.
-However, current RL-based methods still face some limitations.
-Primarily, these methods lack detailed feedback on specific types and distributions of programming errors, which is crucial for identifying patterns, understanding root causes, and creating targeted interventions.
-This makes addressing recurring issues difficult.
-Additionally, the corrective mechanisms are not robust enough, often failing to provide timely, specific guidance, hindering effective learning.
-Additionally, the diversity of input-output examples in existing benchmarks is limited, restricting the model’s ability to adapt and generalize to different or unseen problems [8].
+Reinforcement learning (RL) learns optimal strategies through interaction with the environment and reward signals [14], showing strong performance in sequence generation tasks.
+Code generation, however, is more complex than typical NLP tasks, as it requires not only syntactic correctness but also functional accuracy—ensuring the code compiles and behaves as expected.
+While unit tests are helpful for verifying correctness, they are insufficient unless well-designed to cover diverse and edge cases.
+To address this, methods like CodeRL [6] incorporate unit test feedback into fine-tuning via reinforcement learning.
+PPOCoder [15] improves this with Proximal Policy Optimization, and RLTF [8] further integrates error information and multi-granularity feedback.
+StepCoder [16] enhances the process using compiler feedback and segmental optimization.
+Despite these advances, existing RL-based approaches often lack fine-grained feedback on error types and distributions, limiting their ability to address recurring issues.
+Moreover, their corrective mechanisms are typically weak, offering limited and untimely guidance.
+Finally, the limited diversity in benchmark datasets restricts model generalization to unseen scenarios [8].
 III Problem Setting
-We aim to enhance the automated code generation capabilities of Large Language Models (LLMs) by addressing key challenges in accuracy, diversity, error correction, and code quality.
-Formally, given a high-level specification $D$, the task is to generate a sequence of code tokens $W=\{w_{1},w_{2},\ldots,w_{T}\}$ that maximizes the conditional probability $P(W|D,\theta)$, where $\theta$ represents the model parameters.
-The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\log P(W|D,\theta)\right]$.
-Challenges.
-Effective code generation is impeded by several factors.
-Accuracy requires minimizing syntactical or logical errors in $W$ to ensure correct functionality.
-Diversity of input-output examples $\mathcal{X}$ is often limited, restricting the model’s ability to generalize across varied programming tasks.
-Efficient error correction mechanisms are necessary to identify and rectify errors in $W$, ensuring robust performance.
-Maintaining high code quality, which encompasses adherence to coding standards, style guidelines, and managing code complexity $C(W)$, remains a persistent challenge.
-Additionally, adaptability refers to the model’s capacity to adapt to new tasks and incorporate feedback for continuous improvement, which is constrained without robust memory mechanisms.
-FALCON Framework.
-To address these challenges, we propose the FALCON framework.
-The framework leverages both a long-term memory buffer $\mathcal{M}_{\text{long}}=\{(D_{i},W_{i},T_{i},F_{i})\}_{i=1}^{N}$ and a short-term memory buffer $\mathcal{M}_{\text{short}}=\{(D_{j},W_{j},T_{j},F_{j})\}_{j=1}^{M}$ to utilize diverse feedback, enabling fine-tuning of $\theta$.
-Formally, FALCON seeks to optimize $\theta$ by maximizing a composite reward function $R(W,F)=\alpha T(W)+\beta S(W)+\gamma C(W)+\delta E(W)$, where $T(W)$, $S(W)$, $C(W)$, and $E(W)$ represent unit test, code style, complexity, and error feedback, respectively.
-The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\mathbb{E}_{W\sim P(W|D,\theta)}\left[R(W,F)\right]\right]$.
-Assumptions.
-The effectiveness of FALCON is based on the following assumptions: 1.
-Exchangeability of the dataset $\mathcal{D}=\{D_{i}\}_{i=1}^{N}$, implying that the order of tasks does not influence model performance; 2.
-Independence of feedback signals $F$ given the generated code $W$; 3.
-Sufficient memory capacity in both $\mathcal{M}_{\text{long}}$ and $\mathcal{M}_{\text{short}}$ to store relevant interactions without significant data loss; and 4.
-Feedback efficacy, ensuring that integrated feedback mechanisms provide meaningful and actionable information to guide the optimization of $\theta$.
-Optimization Objective.
-Given a set of tasks $\mathcal{T}=\{D_{i}\}_{i=1}^{N}$, the goal is to learn optimal parameters $\theta^{*}$ that maximize the expected composite reward across all tasks, formulated as $\theta^{*}=\arg\max_{\theta}\frac{1}{N}\sum_{i=1}^{N}R(W_{i},F_{i})$, where $W_{i}\sim P(W|D_{i},\theta)$ is the generated code for task $D_{i}$, and $F_{i}$ represents the aggregated feedback.
-The formulation ensures continuous improvement by leveraging both historical and recent feedback to enhance code generation quality.
+We aim to enhance automated code generation capabilities of Large Language Models (LLMs) by addressing accuracy, diversity, error correction, code quality, and adaptability.
+Formally, given a high-level specification $D$, the task is to generate a sequence of code tokens $W=\{w_{1},w_{2},\ldots,w_{T}\}$ that maximizes $P(W|D,\theta)$, with the objective $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}[\log P(W|D,\theta)]$.
+Accuracy requires minimizing syntactical or logical errors, while limited diversity in input-output pairs $\mathcal{X}$ restricts generalization.
+Efficient error correction is crucial for identifying and fixing errors in $W$.
+Code quality involves adhering to standards, style guidelines, and managing complexity $C(W)$.
+Adaptability depends on the model’s capacity to incorporate feedback for continuous improvement, which is constrained without robust memory.
+The FALCON framework leverages a long-term memory buffer $\mathcal{M}_{\text{long}}=\{(D_{i},W_{i},T_{i},F_{i})\}_{i=1}^{N}$ and a short-term memory buffer $\mathcal{M}_{\text{short}}=\{(D_{j},W_{j},T_{j},F_{j})\}_{j=1}^{M}$ to utilize diverse feedback for fine-tuning $\theta$.
+FALCON optimizes $\theta$ by maximizing a composite reward $R(W,F)=\alpha T(W)+\beta S(W)+\gamma C(W)+\delta E(W)$, where $T(W)$, $S(W)$, $C(W)$, and $E(W)$ denote unit test, style, complexity, and error feedback, respectively, yielding $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}[\mathbb{E}_{W\sim P(W|D,\theta)}[R(W,F)]]$.
+We assume exchangeability of $\mathcal{D}$, independence of feedback $F$ given $W$, sufficient memory capacity in $\mathcal{M}_{\text{long}}$ and $\mathcal{M}_{\text{short}}$, and effective feedback mechanisms.
+Given tasks $\mathcal{T}=\{D_{i}\}_{i=1}^{N}$, the ultimate goal is $\theta^{*}=\arg\max_{\theta}\frac{1}{N}\sum_{i=1}^{N}R(W_{i},F_{i})$ with $W_{i}\sim P(W|D_{i},\theta)$, ensuring continuous improvement by leveraging both historical and recent feedback to enhance code generation quality.
 IV Methodology
-In this section, we explore the FALCON framework, which integrates comprehensive unit testing with reinforcement learning, supported by both long-term and short-term memory buffers.
-During the code generation process, the system stores task descriptions, generated code, and various feedback (e.g., compilation results, code style, and complexity) in the long-term memory buffer.
-By retrieving this information, the model references high-quality code, avoids past mistakes, and ensures adherence to required standards.
-After generating the code, a judge model evaluates it and calculates rewards based on the feedback, which are then used to update the model’s parameters through reinforcement learning.
-All generated code and feedback are stored for future reference and optimization.
-The combination of long-term and short-term memory feedback in the FALCON framework allows the model to not only learn from a wide range of historical data but also adapt quickly to new tasks based on recent performance.
-The overall framework is illustrated in Figure 1.
-Figure 1: The overview of the FALCON framework.
-Detailed feedback process (see Appendix Figures 5 and 6 for feedback cases).
-Reinforcement learning updates using feedback to compute RL loss and optimize the model (see Figure 2 for the detailed meta-reinforcement learning framework).
-IV-A Task Definition: Code Generation by LLMs
-The code generation task involves the automated creation of computer code $W$ from a high-level specification of the desired behavior $D$.
-The goal of this process is to enhance the likelihood of generating code that effectively solves a given problem, which is represented as a conditional probability $P(W|D)$.
-This task can be modeled as a sequence-to-sequence problem, aiming to maximize the conditional probability of generating the correct output given the input and the parameters of the LLM model, mathematically represented as:
-$\max_{\theta}P(W|D,\theta)=\max_{\theta}\prod_{t=1}^{T}p(w_{t}|D,\theta,w_{1:t-1})$ (1)
-where $\theta$ denotes the parameters of the LLM model.
-The LLM model is employed to learn the conditional distribution and is trained on a set of input-output pairs.
-During the training phase, the parameters $\theta$ are optimized to increase the likelihood of generating accurate outputs for a given input.
-IV-B MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
+The FALCON framework enhances code generation by LLMs through unit testing and reinforcement learning, leveraging long-term and short-term memory buffers.
+Task descriptions, generated code, and feedback (e.g., compilation results, code style, complexity) are stored in the long-term memory buffer, allowing the model to reference high-quality code, avoid errors, and adhere to standards.
+A judge model evaluates the code, calculates feedback-based rewards, and updates model parameters via reinforcement learning.
+This dual memory mechanism enables the model to improve code quality using historical data while adapting to new tasks.
+The method section includes three components: Meta-Reinforcement Learning with Dual Memory Buffers IV-A, introducing the dual memory structure; and Long-Term Memory Feedback IV-B, optimizing with historical data.
+See Figure 1.
+IV-A MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
 We propose the MRLF algorithm 1, which involves random task sampling from the task distribution during training.
@@
 $\theta_{\text{final}}=\text{Optimize}\left(\theta_{\text{meta}},\theta,\{\theta_{i}^{\prime}\}\right)$ (5)
-IV-C Long-Term Memory Feedback
-Given that retrieving information from long-term memory helps improve code quality, we use the FAISS framework [22] to retrieve relevant historical code, feedback, and evaluation scores.
-Task descriptions and feedback are transformed into embedding vectors and then indexed.
-During code generation, a query vector from the current task retrieves the top-k most similar historical data to guide the process and avoid past errors.
-The prompt template is provided in the appendix.
-Consider a set of historical data $\mathcal{D}={(t_{i},f_{i},e_{i})}_{i=1}^{n}$, where $t_{i}$ represents the task description, $f_{i}$ is the corresponding feedback, and $e_{i}$ is the evaluation score.
-We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\boldsymbol{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS.
-During the code generation phase, the current task description $\boldsymbol{t}_{\text{current}}$ and feedback $\boldsymbol{f}$ are transformed into a query vector $\boldsymbol{q}=\phi(\boldsymbol{t}_{\text{current}})$.
-We compute the similarity between the query vector $\boldsymbol{q}$ and the historical vectors $\boldsymbol{v}_{i}$ using cosine similarity $\text{cos}(\boldsymbol{q},\boldsymbol{v}_{i})$, and retrieve the top-$k$ most similar historical tasks.
-The retrieval process can be represented as:
-$\{(t_{i_{1}},f_{i_{1}},e_{i_{1}}),\dots,(t_{i_{k}},f_{i_{k}},e_{i_{k}})\}=\text{Top-}k\left(\boldsymbol{v}_{i}\mid i=1,2,\dots,n\right)$ (7)
-By referencing these most relevant historical tasks and feedbacks, the system can guide the current code generation process with past mistakes avoided and ultimate code quality improved.
-IV-D Short-Term Memory Feedback
-During the reinforcement learning phase, we utilize the generated code $\hat{w}$ to construct the reinforcement learning loss function as follows:
-$L_{r1}=-\sum_{t=S_{\text{fine}}}^{E_{\text{fine}}}R_{\text{fine}}(\hat{w}_{t})\log p(\hat{w}_{t}|D,\theta,\hat{w}_{1:t-1})$ (8)
-where $R_{\text{fine}}(\ast)$ represents the reward coefficient, and $S_{\text{fine}}$ and $E_{\text{fine}}$ denote the start and end positions of the code snippet, respectively.
-These values are determined based on different types of feedback.
-Compiler Feedback. For compiler feedback, we adopt the same settings as CodeRL:
-$R_{\text{coarse}}(\hat{W})=\left\{\begin{array}[]{ll}\phantom{-}1.0,&\text{if }FB(\hat{W})\text{ is pass}\\ -0.3,&\text{if }FB(\hat{W})\text{ is failure}\\ -0.6,&\text{if }FB(\hat{W})\text{ is runtime error}\\ -1.0,&\text{if }FB(\hat{W})\text{ is syntax error}\end{array}\right.$ (9)
-where $R_{\text{coarse}}$ is based on compiler feedback with the start position $S_{\text{coarse}}=0$ and end position $E_{\text{coarse}}=T$.
-Adaptive Feedback.
-To enhance the model’s efficiency in handling various programming tasks, we devise a mechanism that dynamically adjusts rewards based on the proportion of passes to failures in unit tests.
-This strategy encourages the model not only to pass unit tests but also to learn from failures, thereby improving its problem-solving capabilities.
-The reward is calculated as:
-$R_{\text{error}}(\hat{W})=-0.3+1.3\times\frac{N_{\text{pass}}}{N_{\text{pass}}+N_{\text{fail}}}$ (10)
-Coding Style Feedback.
-To further enhance the quality of the generated code, we employ AI Feedback to optimize coding style.
-An evaluation model scores the generated code based on adherence to the expected coding style standards.
-The scoring system ranges from -1 to 2, and these evaluation scores are directly used as reward signals in the reinforcement learning process to guide the model toward producing higher-quality code.
-The coding style assessment template is provided in Table X.
-Complexity Feedback. Similar to the coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores. The complexity assessment template is provided in Table XI.
-Long-term Error Type Feedback.
-Introducing a reward mechanism that combines short-term memory recall error rate, current test error, and long-term memory feedback of past task performance enables the model to dynamically adjust its rewards based on past error patterns, adapt to various error types and feedback, and ultimately enhance its generalization ability:
-$R_{\text{negative}}=-\sum_{\text{error}}N_{\text{error}}\times P_{\text{error}}$ (11)
-where $N_{\text{error}}$ represents the frequency of each error type in the generated code, which is the short-term memory feedback and the immediate feedback for the current task.
-$P_{\text{error}}$ represents the proportion of each error type, which is the long-term memory feedback.
-This is the model’s performance statistics on long-term tasks, providing a cumulative history of various error types.
-By correct rewarding, the model can reduce the occurrence of these errors and thereby enhance the accuracy and quality of the generated code.
 TABLE I: Quantitative evaluation on the APPS benchmark.
@@
 Ours 770M 8.60 2.56 1.24 3.50 19.75 5.85 3.57 8.17
+IV-B Long-Term Memory Feedback
+Retrieving information from long-term memory helps improve code quality.
+We use the FAISS framework [19] to retrieve relevant historical code, feedback, and evaluation scores.
+Task descriptions and feedback are transformed into embedding vectors and then indexed.
+During code generation, a query vector from the current task retrieves the top-k most similar historical data to guide the process and avoid past errors.
+The prompt template is provided in the appendix.
+Consider a set of historical data $\mathcal{D}={(t_{i},f_{i},e_{i})}_{i=1}^{n}$, where $t_{i}$ represents the task description, $f_{i}$ is the corresponding feedback, and $e_{i}$ is the evaluation score.
+We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\boldsymbol{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS.
+During the code generation phase, the current task description $\boldsymbol{t}_{\text{current}}$ and feedback $\boldsymbol{f}$ are transformed into a query vector $\boldsymbol{q}=\phi(\boldsymbol{t}_{\text{current}})$.
+We compute the similarity between the query vector $\boldsymbol{q}$ and the historical vectors $\boldsymbol{v}_{i}$ using cosine similarity $\text{cos}(\boldsymbol{q},\boldsymbol{v}_{i})$, and retrieve the top-$k$ most similar historical tasks.
+The retrieval process can be represented as:
+$\{(t_{i_{1}},f_{i_{1}},e_{i_{1}}),\dots,(t_{i_{k}},f_{i_{k}},e_{i_{k}})\}=\text{Top-}k\left(\boldsymbol{v}_{i}\mid i=1,2,\dots,n\right)$ (6)
+By referencing these most relevant historical tasks and feedbacks, the system can guide the current code generation process with past mistakes avoided and ultimate code quality improved.
+IV-C Short-Term Memory Feedback
+During the reinforcement learning phase, we utilize the generated code $\hat{w}$ to optimize the policy based on the reinforcement learning loss defined in Equation (1).
+As previously described, this loss incorporates reward signals such as unit test feedback, code complexity, and style norms, and is computed over the token span from $S_{\text{fine}}$ to $E_{\text{fine}}$.
+Compiler Feedback. For compiler feedback, we adopt the same settings as CodeRL:
+$R_{\text{coarse}}(\hat{W})=\left\{\begin{array}[]{ll}\phantom{-}1.0,&\text{if }FB(\hat{W})\text{ is pass}\\ -0.3,&\text{if }FB(\hat{W})\text{ is failure}\\ -0.6,&\text{if }FB(\hat{W})\text{ is runtime error}\\ -1.0,&\text{if }FB(\hat{W})\text{ is syntax error}\end{array}\right.$ (7)
+$S_{\text{coarse}}=0,\quad E_{\text{coarse}}=T$
+where $R_{\text{coarse}}$ is based on compiler feedback with the start and end positions set to 0 and $T$.
+Adaptive Feedback.
+To enhance the model’s efficiency in handling various programming tasks, we devise a mechanism that dynamically adjusts rewards based on the proportion of passes to failures in unit tests.
+This strategy encourages the model not only to pass unit tests but also to learn from failures, thereby improving its problem-solving capabilities.
+The reward is calculated as:
+$R_{\text{error}}(\hat{W})=-0.3+1.3\times\frac{N_{\text{pass}}}{N_{\text{pass}}+N_{\text{fail}}}$ (8)
+Coding Style Feedback.
+To further enhance the quality of the generated code, we employ AI Feedback to optimize coding style.
+An evaluation model scores the generated code based on adherence to the expected coding style standards.
+The scoring system ranges from -1 to 2, and these evaluation scores are directly used as reward signals in the reinforcement learning process to guide the model toward producing higher-quality code.
+Complexity Feedback. Just like with coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores.
+Long-term Error Type Feedback. A reward mechanism combining short-term error recall, current test errors, and long-term memory of past performance dynamically adjusts rewards. The formula is:
+$R_{\text{negative}}=-\sum_{\text{error}}N_{\text{error}}\times P_{\text{error}}$ (9)
+where $N_{\text{error}}$ represents short-term feedback on current task errors, and $P_{\text{error}}$ represents long-term error proportions.
+This mechanism reduces errors and enhances the accuracy and quality of code generation.
 V Experiment
@@
 V-B Quantitative Evaluation on HumanEval and MBPP
-To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model trained with our method on APPS, using the well-established MBPP and HumanEval benchmarks.
+To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model, trained with our method on our custom dataset, using the well-established MBPP and HumanEval benchmarks.
 We also compare these results against other reinforcement learning methods, such as PPOCoder and RLTF.
@@
 This approach demonstrates a strong ability to enhance the generation of correct code and proves particularly effective in complex tasks.
-TABLE III: Test set results of AGENTBENCH.
-Model Size VER OS DB KG
-GPT-3.5-turbo - 0613 31.6 15.7 25.9
-GPT-4 - 0613 42.4 32.0 58.8
-Tinyllama 1.1B - 2.8 0 0
-Codellama 7B instruct 9.7 2.7 0
-Qwen 7B chat 12.5 13.0 7.0
-Agentlm 7B chat 14.6 33 9.0
-Deepseek-Coder 6.7B instruct 17.4 23.3 6.8
-DeepSeek-Coder (FALCON) 6.7B instruct 22.2 26.7 9.0
 V-C Quantitative Evaluation on CODAL-Bench
@@
 Figure 3: Quantitative evaluation on CODAL-Bench
-V-D Quantitative Evaluation on SciCode
-To validate the general-purpose task capabilities of our framework, we also select the SciCode benchmark, which covers challenging research-level coding problems across natural sciences, including mathematics, physics, chemistry, and biology.
-SciCode decomposes the main problems into several subproblems, making it a particularly rigorous benchmark of a model’s coding capabilities.
-Even the most advanced models nowadays, such as Claude 3.5-Sonnet and ChatGPT-4.0, can only solve 1.5% and 4.6% of the main problems, respectively.
-Although Deepseek-Coder-6.7B-instruct initially demonstrates a low task pass rate on this benchmark, we observe improvements on the subproblems after applying our framework due to the utilization of long-term memory mechanisms.
-TABLE IV: The results of pass@1 on SciCode benchmarks with and without FALCON.
-Model Size Method Subproblem Main Problem
-CodeLlama 70B - 10.4 0
-CodeLlama 7B w/o 0.4 0
-CodeLlama 7B w 3.5 0
-DeepSeek-Coder 6.7B w/o 5.2 0
-DeepSeek-Coder 6.7B w 8.3 0
-V-E Quantitative Evaluation on Agentbench
-To further evaluate the performance of our framework, we conduct a systematic assessment on AgentBench [23], focusing specifically on long-term memory capabilities.
-Since our primary focus is on code generation tasks, we select three environments within AgentBench: Operating System (OS), Database (DB), and Knowledge Graph (KG).
-In this evaluation, we compare proprietary models (such as GPT-4 and GPT-3.5) with open-source models (such as Codellama and Qwen 2.5) [24].
-The results reveal that the models optimized through our framework exhibit significant improvements, particularly in the OS environment with an increase of 4.8 percentage points.
-The experimental results are illustrated in Table III.
-V-F Ablation Studies
+V-D Ablation Studies
 The Influence of Models.
-To validate the scalability and robustness of our framework, we conduct experiments with DeepSeek-Coder-Instruct-6.7B.
+To validate the scalability and robustness of our framework, we conduct experiments with the larger model, DeepSeek-Coder-Instruct-6.7B, to further evaluate its performance.
 Notably, the improvements in introductory-level tasks are significant, which can be attributed to the use of long-term memory that enhances the quality of generated data and further unlocks the model’s potential.
-The results are illustrated in Table V.
-TABLE V: Different large language models as the backbone
+The results are illustrated in Table III.
+TABLE III: Different large language models as the backbone
 Model Size Method Intro Inter Comp All
@@
 DeepSeek-Coder 6.7B w 22.40 8.52 3.70 10.33
-The Influence of Different Feedbacks on Coding Preferences.
-As shown in Figure 4, ablation experiments are also conducted to validate the effectiveness of the feedback that we introduce for coding preferences.
-It is found that incorporating targeted feedback enhances the model’s performance concerning its respective coding preferences.
-Notably, the optimization aiming at increasing code complexity achieves the best results.
-Although there are some improvements in coding style and instruction following, it is worth noting that the enhancement in code instruction following is not significant.
-We leave it as a topic for future investigation.
-Figure 4: Effect of different feedbacks on coding preferences
 The Influence of Memory.
 To validate the impact of long-term and short-term memories on code generation capabilities, we conduct ablation experiments using CodeT5 as the base model and test it on the APPS dataset.
-As shown in Table VI, the experimental results indicate that both long- and short-term memories enhance the model’s code generation performance effectively, while the short-term memory feedback demonstrates a more significant improvement attributed to the reward design which plays an important role in fine-tuning the model.
-TABLE VI: Effect of long and short memories on different performance metrics
+As shown in Table IV, the experimental results indicate that both long- and short-term memory feedback enhances the model’s code generation performance effectively, while the short-term memory feedback demonstrates a more significant improvement.
+This improvement can be attributed to the effective reward design which plays a positive role in fine-tuning the model.
+TABLE IV: Effect of long and short memories on different performance metrics
 Long Memory Short Memory Intro Inter Comp All
@@
 8.60 2.56 1.25 3.50
-VI Conclusions and Future Work
-In this work, we propose FALCON, a novel RL-based feedback framework that enhances automatic code generation by integrating long-term and short-term memory feedback within a meta-reinforcement learning strategy to better capture human intent like RL based preference learning.
+VI Conclusions
+In this work, we propose FALCON, a novel RL-based feedback [20] framework that enhances automated code generation by integrating long-term and short-term memory feedbacks within a meta-reinforcement learning strategy to better capture human intent like RL based preference learning.
 Long-term memory retains past interactions to improve code quality and reduce repetitive mistakes, while short-term memory enables immediate adjustments based on recent feedback from compilers and AI systems.
@@
 Extensive evaluations on benchmarks including APPS, HumanEval, and CODAL-Bench demonstrate that FALCON outperforms existing RL-based methods, achieving higher functional correctness and better coding style adherence.
-In future work, we aim to expand FALCON’s capabilities by incorporating a broader diversity of programming languages and tackling more complex code generation challenges.
-Appendix A Example
-We provided an example of code generation with long-term memory.
-When generating code without long-term memory, it often results in repetitive ValueError issues.
-By incorporating long-term memory to retrieve the most relevant code blocks and embedding them as context during generation, the quality of the generated code can be significantly improved.
-We also provided an instance of using long-term memory in SCICode.
-In direct generation, there were certain logical issues within the code, and the coding style lumped all formulas together.
-However, after employing long-term memory retrieval for assistance, the code was segmented appropriately and the logic was correctly implemented.
-Figure 5: An example of code generation with long-term memory incorporated.
-Figure 6: An example of code generation with long-term memory incorporated.
-Appendix B Feedback Error Correlation Analysis
-We performed an analysis of different feedback types and their associated error categories.
-We conducted experiments using CodeT5 as the base model on APPS with individual feedback and collected the occurrence frequency of various sub-errors.
-The results are shown in the Figure 7.
-The experimental results indicate that compiler feedback significantly reduces Syntax Errors and Index Errors.
-However, it also slightly increases the occurrence of Value Errors.
-This can be attributed to the corrective nature of compiler feedback on errors.
-Other feedback types, such as Coding Style, Instruction Following, and Code Complexity Feedback, can reduce Syntax Errors compared to having no feedback.
-However, their reduction is not as significant as that achieved by compiler feedback.
-Instruction Following Feedback specifically shows some improvement in reducing Value Errors, indicating an enhancement in instruction adherence.
-Figure 7: Frequency of sub-errors generated by different feedback types.
-Appendix C Prompts
-We have compiled relevant prompt templates for code generation based on long-term memory retrieval and AI feedback.
-TABLE VII: Code generation template
-Instruction:
-Please write a Python function based on the task description, referencing the historical context for inspiration.
-Ensure that the generated code follows the provided requirements and avoids the listed errors.
-Task: [instruction]
-Context of relevant code:
-- Historical Task: [Brief description of a similar task]
-- Code: [Code snippet]
-- Style Score: [Style score]
-- Efficiency Score: [Efficiency score]
-- Additional Feedback: [Additional comments or issues]
-Requirements:
-1. Ensure the generated code adheres to best practices for Python, including proper naming conventions, consistent formatting, and coding standards.
-2. Optimize the code for performance, avoiding unnecessary recursion or nested loops.
-3. Use built-in or efficient library functions whenever applicable to improve both readability and performance.
-Avoid the following errors:
-- [Historical error] – Avoid structural or logical issues found in previous code snippets.
-Output: Ensure your response is in the format of ”’python”’.
-TABLE VIII: Results of pass@1 on APPS.
-Models Size Introductory Interview Competition Overall
-Base Models
-CodeLlama 13B 14.0 5.0 1.0 6.0
-CodeLlama-Python 13B 25.3 8.2 2.4 10.5
-DeepSeek-Coder-Base 6.7B 11.4 8.5 2.8 7.9
-Supervised Fine-tuned Models
-StarCoder 15.6B 6.2 2.9 0.3 3.1
-CodeLlama-Instruct 13B 30.1 11.8 3.2 13.7
-WizardCoder-Python-V1.0 13B 35.2 10.9 4.1 14.4
-DeepSeek-Coder-Instruct 6.7B 40.2 13.7 2.7 16.8
-Reinforcement Learning-based Models (DeepSeek-Coder-Instruct-6.7B as the backbone)
-PPOCoder 6.7B 41.2 14.0 2.2 17.1
-CodeRL 6.7B 40.8 13.7 2.7 16.9
-RLTF 6.7B 42.4 13.3 3.1 17.1
-FALCON(Ours) 6.7B 44.7 14.4 3.2 18.2
-TABLE IX: Results of pass@1 on LeetCode.
-Models Size Easy(45) Medium(91) Hard(44) Overall(180)
-GPT-3.5-Turbo - 46.7 15.4 15.9 23.3
-WizardCoder-Python-V1.0 13B 17.8 1.1 0 5
-CodeLlama-Instruct 34B 24.4 4.4 4.5 9.4
-DeepSeek-Coder-Instruct 6.7B 44.4 12.1 9.1 19.4
-Reinforcement Learning-based Models
-PPOCoder 6.7B 48.9 13.1 9.1 21.1
-CodeRL 6.7B 55.6 14.2 9.1 23.3
-RLTF 6.7B 57.8 13.1 11.3 23.9
-FALCON(Ours) 6.7B 60.0 15.4 11.3 25.6
-TABLE X: Coding style assessment template.
-Coding Style Assessment
-Evaluate the coding style of provided code segments.
-Assess how well the code adheres to the best practices of the language, focusing on readability, maintainability, and efficiency in line with the language’s idiomatic style.
-Reward Scale: Rate outputs on a scale of -1 to 2
--1. Poor Adherence: The code significantly deviates from standard practices, showing poor readability, maintainability, and efficiency.
-0. Basic Adherence: The code makes some effort to follow language conventions but lacks consistency in readability, maintainability, or efficiency.
-1. Good Adherence: The code generally follows standards, demonstrating adequate readability, maintainability, and efficiency, though with room for improvement.
-2. Excellent Adherence: The code exemplifies best practices, with high readability, maintainability, and efficiency, fully adhering to idiomatic conventions.
-TABLE XI: Complexity assessment template.
-Complexity Assessment
-Evaluate the solutions and code provided by the assistant based on their complexity. Assess how well the code manages complexity while achieving the desired outcomes.
-Reward Scale: Rate outputs on a scale of -1 to 2
--1. Overly Complex: The code is unnecessarily complicated, with a high level of complexity that makes it hard to understand or maintain.
-0. Acceptable Complexity: The code has a reasonable level of complexity, but there may be opportunities to simplify further.
-1. Moderately Simple: The code is simple and well-organized, with minimal complexity and clear logic.
-2. Optimal Simplicity: The code exemplifies the best practices in minimizing complexity while ensuring functionality.
-TABLE XII: Instruction following assessment template.
-Instruction Following Assessment
-Evaluate the assistant’s fidelity to provided instructions. Assess how accurately the assistant’s responses align with user directives, noting any deviations and their justification.
-Evaluation Criteria
-Precision in Following Instructions: Does the assistant adhere to the specifics of the provided instructions?
-Justification for Deviations: If deviations occur, are they justified by critical necessity or explicit user request?
-Alignment with User Directives: How well do the assistant’s responses match the user’s specified needs and expectations?
-Necessity of Deviations: Are any deviations from instructions made only in situations deemed absolutely necessary or upon direct user request?
-Reward Scale: Rate outputs on a scale of -1 to 2
--1. Non-Compliant: The assistant frequently deviates from instructions without necessity or user consent.
-0. Acceptable: The assistant shows some adherence to instructions but deviates without strong justification.
-1. Compliant with Justified Deviations: The assistant generally follows instructions, with deviations occurring but justified by necessity or user request.
-2. Fully Compliant: The assistant follows instructions closely, with minimal deviations, all of which are well justified.
-Appendix D Error Category
-Due to the differences in languages accepted by Compiler Feedback during unit tests for various language tasks, we have standardized the definition of sub-errors in Compiler Feedback.
-The table XIII XIV XV below outlines our specifications for Python, C, and Java.
-TABLE XIII: Common Python errors with categories.
-Sub-error Description Category
-Syntax Error Code contains syntax errors that cause the compilation to fail. Syntax Error
-Indentation Error Wrong indentation format. Syntax Error
-Index Error Index operation is out of bounds. Runtime Error
-Type Error An operation or function was applied to an object of an inappropriate type. Runtime Error
-Value Error An operation or function received an argument with the correct type but with an inappropriate value. Runtime Error
-EOF Error The input() function encountered an end-of-file condition (EOF) without reading any data. Runtime Error
-Timeout Error Code execution time exceeds time limit. Runtime Error
-Name Error A local or global name is not defined. Runtime Error
-Key Error A mapping (dictionary) key is not found in the set of existing keys. Runtime Error
-Import Error The imported package is not found. Runtime Error
-ZeroDivision Error The second argument of a division or modulo operation is zero. Runtime Error
-Recursion Error Code execution recursive operation exceeds the maximum limit. Runtime Error
-TABLE XIV: Common C Language errors with categories.
-Sub-error Description Category
-Segmentation Fault Accessing memory that the program doesn’t have permission to access. Runtime Error
-Null Pointer Dereference Attempting to dereference a pointer that is NULL. Runtime Error
-Buffer Overflow Writing data outside the allocated buffer memory. Runtime Error
-Memory Leak Dynamically allocated memory not being freed. Runtime Error
-Syntax Error A syntax mistake in the code, such as a missing semicolon. Syntax Error
-Type Mismatch Assigning a value of one type to a variable of another type. Syntax Error
-Uninitialized Variable Using a variable before it has been initialized. Runtime Error
-Undefined Behavior Code that can exhibit unpredictable behavior depending on compiler or runtime environment. Runtime Error
-Division by Zero Attempting to divide a number by zero. Runtime Error
-Infinite Loop A loop that never terminates due to incorrect logic. Runtime Error
-TABLE XV: Common Java Language errors with categories.
-Sub-error Description Category
-NullPointerException Attempting to access an object with a null reference. Runtime Error
-ArrayIndexOutOfBounds- Exception Accessing an array index that is out of bounds. Runtime Error
-ClassCastException Casting an object to a subclass it is not an instance of. Runtime Error
-NumberFormatException Attempting to convert a string to a number, but the string doesn’t have the appropriate format. Runtime Error
-StackOverflowError Recursive method calls that exceed the stack size. Runtime Error
-Syntax Error Any mistake in the code structure such as missing braces or semicolons. Syntax Error
-ClassNotFoundException The Java class is not found at runtime. Runtime Error
-IllegalArgumentException A method has been passed an illegal or inappropriate argument. Runtime Error
-ArithmeticException Division by zero or other illegal arithmetic operations. Runtime Error
-UnsupportedOperation- Exception When a requested operation is not supported. Runtime Error
-Appendix E Supplementary Experiments
-E-A Quantitative Evaluation on APPS
-We utilized a larger model, DeepSeek-Coder-Instruct-6.7B, as the backbone to compare popular and widely used advanced reinforcement learning methods, further validating the effectiveness and necessity of our approach.
-The specific results indicate that our method achieved scores of 44.7, 14.4, and 3.2 in the categories of "Introductory" "Interview" and "Competition" respectively.
-This demonstrates our method’s exceptional capability in exploring the output space of larger models, highlighting its versatility.
-E-B Quantitative Evaluation on LeetCode
-To further validate our approach in solving real-world programming problems, we utilized the LeetCode Contest benchmark.
-This benchmark comprises 180 of the latest LeetCode contest problems collected from July 2023 to January 2024.
-Each problem is accompanied by 100 test cases, and the data format is consistent with human-eval.
-The results indicate that our approach can enhance the code generation capabilities of models in solving real-world programming problems.
-Compared to other RL-based methods, our approach achieved greater improvements and can rival models with significantly larger parameters.